### v1

# Exploring Imbalanced Annotations for Effective In-Context Learning

Abstract
Large language models (LLMs) have shown impressive performance on downstream tasks through in-context learning (ICL), which heavily relies on the demonstrations selected from annotated datasets. Existing selection methods may hinge on the distribution of annotated datasets, which can often be long-tailed in real-world scenarios. In this work, we show that imbalanced class distributions in annotated datasets significantly degrade the performance of ICL across various tasks and selection methods. Moreover, traditional rebalance methods fail to ameliorate the issue of class imbalance in ICL. Our method is motivated by decomposing the distributional differences between annotated and test datasets into two-component weights: class-wise weights and conditional bias. The key idea behind our method is to estimate the conditional bias by minimizing the empirical error on a balanced validation dataset and to employ the two-component weights to modify the original scoring functions during selection. Our approach can prevent selecting too many demonstrations from a single class while preserving the effectiveness of the original selection methods. Extensive experiments demonstrate the effectiveness of our method, improving the average accuracy by up to 5.46 on common benchmarks with imbalanced datasets.
Machine Learning, ICML
1 Introduction
In-context learning (ICL) has been a remarkable capability of Large Language Models (LLMs), enabling them to perform downstream tasks by simply conditioning on a few task demonstrations. (Brown et al., 2020). ICL consistently outperforms zero-shot inference across various tasks without needing parameter updates, positioning it as a strong alternative to supervised fine-tuning (SFT) (Mosbach et al., 2023; Panwar et al., 2024). In particular, previous studies typically focus on designing methods for demonstration selection from high-quality annotated datasets (Liu et al., 2022; Baldassini et al., 2024; Wang et al., 2024). The effectiveness of these selection algorithms may hinge on the distribution of annotated datasets, which can often be long-tailed in real-world scenarios (Wei et al., 2022; De Alvis & Seneviratne, 2024) — certain classes are underrepresented (see Figure 1). Thus, it is of great importance to investigate how such imbalances affect ICL’s performance.
Figure 1: An example of imbalanced dataset. In Emotion (Saravia et al., 2018), a few sentiments make a large contribution and data tend to show a long-tailed distribution. For example, “joy“ and “sadness“ are head classes, while most other classes, such as “love“ and “surprise“, are tail classes.
In this work, we present the first study on the impact of long-tailed distribution in annotated datasets on in-context learning. First, we theoretically show that the class prior distribution affects the prediction of ICL. Empirically, we also find that imbalanced class distributions in annotated datasets significantly degrade the performance of ICL across various tasks and selection methods (refer to Figure LABEL:figure:2). Moreover, traditional rebalance methods not only fail to ameliorate but, in some cases, exacerbate the issue of class imbalance in ICL. This motivates us to develop a universal method that consistently improves ICL’s performance in the presence of imbalanced annotated datasets.
In this paper, we show that the issue of imbalanced annotated datasets can be mitigated by applying class weighting to the original scoring functions during selection. Our method is motivated by accounting for the differences between annotated and test distributions using two-component weights: class-wise weights, which assesses the difference between class prior distributions, and conditional bias, which measures the bias of class-wise weights due to different conditional distributions. Therefore, our key idea is to employ the two-component weights to rectify the original scoring functions during selection. Specifically, we employ effective number (Cui et al., 2019) as class-wise weights and estimate conditional bias by minimizing the empirical error of a balanced validation set. We then employ the two-component weights to modify the original scoring functions during selection. In effect, our approach can prevent selecting too many demonstrations from a single class while preserving the effectiveness of the original selection methods.
[…]
Our contributions are summarized as follows:
  1. 1.
    We find that imbalanced class distributions in annotated datasets significantly degrade the ICL’s performance. We also show that classical rebalance methods fail to ameliorate the imbalanced problems of ICL.
  2. 2.
    We propose a simple and effective method to enhance ICL’s performance with imbalanced annotated datasets. The key idea is to estimate the conditional bias and employ the two-component weights to modify the original scoring functions during selection.
  3. 3.
    We empirically show that our method can improve the performance of existing selection methods across various imbalance ratios. Meanwhile, we also validate the effectiveness of our method on generation tasks.
2 Preliminary
2.1 In-context learning
In the context of large language models (LLMs), in-context learning (ICL) aims to generate text outputs $\mathrm{\mathbf{y}}=(y_{1},...,y_{|\mathrm{\mathbf{y}}|})$ (i.e., token sequences) conditioned on input $\mathrm{\mathbf{x}}=(x_{1},...,x_{|\mathrm{\mathbf{x}}|})$ and context $\mathrm{\mathbf{C}}_{K}$. In particular, the context $\mathrm{\mathbf{C}}_{K}=\{(\mathrm{\mathbf{x}}_{i},\mathrm{\mathbf{y}}_{i})\}^{K}_{i=1}$ comprises $K$ task demonstrations (e.g. input-output pairs) selected from a large annotated dataset with $N$ examples $\mathcal{D}_{c}=\{(\mathrm{\mathbf{x}}_{i},\mathrm{\mathbf{y}}_{i})\}^{N}_{i=1}$. Let $f_{\theta}(\mathrm{\mathbf{C}}_{K},\cdot)$ be the ICL model with demonstrations $\mathrm{\mathbf{C}}_{K}$, using the LLM $f$ parameterized by $\boldsymbol{\theta}$. Given a test input $\mathrm{\mathbf{x}}_{t}$, we generate the output $\mathrm{\mathbf{y}}_{t}$ via the ICL model as:
$\mathrm{\mathbf{y}}_{t}=f_{\theta}(\mathrm{\mathbf{C}}_{K},\mathrm{\mathbf{x}}_{t})=\mathop{\arg\max}\limits_{\mathrm{\mathbf{y}}}P_{\theta}\left(\mathrm{\mathbf{y}}_{t}|\mathrm{\mathbf{C}}_{K},\mathrm{\mathbf{x}}_{t}\right).$ (1)
To improve the performance of ICL, previous studies (Liu et al., 2022; Peng et al., 2024; Rubin et al., 2022; Ye et al., 2023a) designed various scoring functions $s(\cdot,\cdot)$ to select demonstrations $\mathrm{\mathbf{C}}_{K}$ from an annotated dataset $\mathcal{D}_{c}$ as:
$\mathrm{\mathbf{C}}_{K}=\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right).$ (2)
[…]
While current selection methods showcase promising performance on common-used benchmarks, their effectiveness may hinge on the distribution of annotation datasets $\mathcal{D}_{c}$. For example, ICL might struggle to make accurate predictions for underrepresented groups within these annotated datasets. Real-world datasets (see Figure 1), however, often exhibit an imbalanced distribution, with a few ‘head’ classes containing many examples and numerous ‘tail’ classes having significantly fewer examples. The concern may lead to challenges in effectively employing in-context learning in real-world applications. We proceed with a formulation of the imbalanced setting of ICL.
2.2 Setting of imbalanced ICL
In this work, we begin with the class-imbalance setting of in-context learning in classification tasks [^1] In addition to labels, the imbalance can also occur among various groups, particularly in generation tasks. We extend the imbalanced setting to generation tasks in Section 6. , where the label space $\mathcal{Y}:=\{1,\ldots,K\}$. Let $n_{j}$ denote the number of instances in class $j$, where $j\in\mathcal{Y}$. In the class-imbalanced setting, the annotated dataset $\mathcal{D}_{c}$ has an unequal distribution of instances across different classes in $\mathcal{Y}$. In particular, the class distribution is such that: $n_{j}\ll n_{K}$, for some $j,k\in\mathcal{Y}$, where $j\neq K$. We quantify the imbalance ratio as $\phi=\frac{\min_{j\in\mathcal{Y}}n_{j}}{\max_{j\in\mathcal{Y}}n_{j}}$, and a higher imbalance ratio indicates a more severe class imbalance in the dataset.
In the real world, class-imbalanced distributions are frequently observed in various datasets. For instance, in the Emotion dataset (Saravia et al., 2018), the ‘Joy’ class constitutes 33% of the data, whereas the ‘Surprise’ class makes up only 4%. The CodeSearchNet dataset (Husain et al., 2019) includes 29% JavaScript code while Ruby accounts for just 3%, demonstrating the significant imbalance issue among programming languages. Therefore, it is crucial to ensure the ICL performance across all classes under the class-imbalanced setting in $\mathcal{D}_{c}$. In what follows, we analyze the effect of class imbalance on ICL from theoretical and empirical perspectives.
3 Pilot study
3.1 The effect of imbalanced dataset
3.1.1 Theoretical Analysis
We start by presenting Remark 3.1 (Xie et al., 2022), which demonstrates that ICL allows LLMs $f$ parameterized by $\theta$ to learn the context-generated distribution from demonstrations $\mathrm{\mathbf{C}}_{K}$, such that the model’s prediction approximates the output $\mathrm{\mathbf{y}}$ generated by the data generated model $\theta^{*}$ given the test input $\mathrm{\mathbf{x}}$.
Remark 3.1.
Assume both demonstrations $\mathrm{\mathbf{C}}_{K}$ and test sample $(\mathrm{\mathbf{x}}_{t},\mathrm{\mathbf{y}}_{t})$ are generated by data generated model $\theta^{*}$. Given such demonstrations $\mathrm{\mathbf{C}}_{K}$, in-context learning allows large language models to generate output $\mathrm{\mathbf{y}}$ as follows:
$f_{\theta}(\mathrm{\mathbf{C}}_{K},\mathrm{\mathbf{x}}_{t})\approx\mathop{\arg\max}\limits_{\mathrm{\mathbf{y}}}P_{\theta^{*}}(\mathrm{\mathbf{y}}_{t}|\mathrm{\mathbf{x}}_{t}).$
where large language models $\theta$ to generate correct output $\mathrm{\mathbf{y}}_{t}$ following context-generated distribution.
From a Bayesian perspective, the prediction of ICL is generally made as follows:
$f_{\theta}(\mathrm{\mathbf{C}}_{K},\mathrm{\mathbf{x}}_{t})\approx\mathop{\arg\max}\limits_{\mathrm{\mathbf{y}}}P_{\theta^{*}}(\mathrm{\mathbf{x}}_{t}|\mathrm{\mathbf{y}}_{t})P_{\theta^{*}}(\mathrm{\mathbf{y}}_{t}).$ (3)
where $P_{\theta^{*}}(\mathrm{\mathbf{x}}_{c})$ and $P_{\theta^{*}}(\mathrm{\mathbf{y}}_{t})$ represent $P_{\theta^{*}}(X=\mathrm{\mathbf{x}}_{t})$ and $P_{\theta^{*}}(Y=\mathrm{\mathbf{y}}_{t})$, respectively.
The Eq. 3 indicates that the class prior distribution $P(Y)$ affects the prediction of ICL. in real-world tasks, the class prior of the test distribution is usually balanced (i.e., a uniform distribution over labels), while the annotated dataset exhibits a long-tailed class distribution. Formally, we have $P_{c}(Y=j)\neq P_{t}(Y=j)$ for any class $j\in\mathcal{Y}$. The predicted posterior probability in Eq. 3 becomes unreliable when there is a large discrepancy of class priors between annotated dataset $P_{c}(Y)$ and test dataset $P_{t}(Y)$.
3.1.2 Empirical Analysis
We also conduct experiments on various downstream tasks, including Amazon (Keung et al., 2020), AgNews, Yahoo and Yelp (Zhang et al., 2015), NQ (Kwiatkowski et al., 2019) and CodeSearchNet (Husain et al., 2019)). To simulate the class-imbalanced setting, we generate imbalanced datasets with a pre-defined probability (e.g., $\phi$ = 0, 10, 50, 100). We evaluate the performance of ICL on a balanced test dataset with three LLMs, including Llama3-8B (AI@Meta, 2024), Mistral-7B (Jiang et al., 2023) and OPT-6.7B (Zhang et al., 2022a). Furthermore, we also compare the performance of ICL with various set sizes and selection methods, including Random (Min et al., 2022), TopK (Liu et al., 2022), DPP (Ye et al., 2023a), VoteK (Su et al., 2023), ConE (Peng et al., 2024), ByCS (Wang et al., 2024).
[…]
Through the above analysis, we find that imbalanced datasets significantly hurt the ICL’s performance. Next, we will verify whether employing some classical rebalance methods can mitigate the issue of imbalanced datasets.
3.2 The failure of classical rebalance methods
Rebalance methods are widely used to deal with the issue of imbalanced annotated datasets. One such approach is to address the bias towards the minor class by adjusting the annotated dataset. In this section, we use four different rebalance methods, including over-sampling (Chawla et al., 2002), under-sampling (Liu et al., 2008), stratified sampling (Vilariño et al., 2005) and re-weighting (Cui et al., 2019). Specifically, given an annotated dataset with $N$ examples, we select demonstrations as follows:
$\mathrm{\mathbf{C}}_{K}$ $=\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}^{{}^{\prime}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N^{{}^{\prime}}}\right),$ (4)
$\mathrm{\mathbf{C}}_{K}$ $=\cup_{i=1}^{k}\operatorname{top}_{\frac{K}{k}}\left(\left\{s(\mathrm{\mathbf{c}}_{j},\mathrm{\mathbf{x}}_{t})\right\}_{j=1}^{n_{i}}\right),$ (5)
$\mathrm{\mathbf{C}}_{K}$ $=\operatorname{Top}_{K}\left(\left\{w_{i}\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right).$ (6)
For the over-sampling method in Eq. 4, we select demonstrations $\mathbf{c}^{{}^{\prime}}_{i}$ from an over-sampling dataset with $N^{{}^{\prime}}$ examples where we repeat the examples of tail classes until their number matches that of head classes. The under-sampling method in Eq. 4 randomly trims examples belonging to head classes until the number of examples in head classes is equal to that of tail classes. Suppose $n_{j}$ represents the size of the class to which $i$-th example belongs, stratified sampling in Eq. 5 selects $\frac{K}{k}$ demonstrations from each class with $n_{j}$ examples. For the re-weighting method in Eq. 6, according to the classical reweighing method for fine-tuning (Cui et al., 2019), we select the top $K$ examples based on the scoring functions $s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})$ multiplied by the class-wise weights $w_{j}$ where $w_{j}=(1-\alpha)/(1-\alpha^{n_{j}})$ and $\alpha=(N-1)/N$.
We conduct experiments to verify the effectiveness of these rebalance strategies across various downstream tasks within the framework of ICL. The results in Figure 3 show that employing over-sampling and under-sampling can improve ICL’s performance when annotated datasets exhibit long-tail distributions. However, these rebalance strategies can only achieve limited effects for most downstream tasks, indicating poor generalization to ICL. Specifically, the limited effect of the oversampling method (Chawla et al., 2002) suggests that repeated examples do not provide additional information to LLMs to enhance the performance of ICL. Under-sampling (Liu et al., 2008) may remove key information of head classes. Notably, the failure of the stratified sampling method in ICL is mainly due to the decreased performance of selection methods such as TopK (Liu et al., 2022) and DPP (Ye et al., 2023a). The detailed results of generation tasks have been reported in Section 6.1.
Our empirical analysis shows that existing rebalance methods cannot effectively address the issue of imbalanced datasets for ICL. This motivates us to design new rebalance methods for ICL that can universally enhance the performance of existing selection methods.
Figure 3: Comparison of performances on four classification tasks with different imbalance ratios using four different rebalance methods. We report the average accuracy (%) of six existing selection methods ($K=8$) using OPT-6.7B (Zhang et al., 2022a).
4 Methodology
In this section, we first decompose the distributional differences between annotated and test datasets to explain why classical rebalance methods fail for ICL. In light of this, we propose a novel method to enhance the performance of ICL with imbalanced annotated datasets.
4.1 Decomposition of the distributional differences
We aim to design a suitable demonstration selection method, which is used to select demonstrations for the given test input $\mathrm{\mathbf{x}}_{t}$, to minimize expected $\operatorname{Error}$ value across all classes:
$\operatorname{Error}=\mathbb{E}_{P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right],$
where $M[\cdot,\cdot]$ measures the mismatched degree between real output and the generated output conditioned on demonstrations $\mathrm{\mathbf{C}}_{K}$ and test input $\mathrm{\mathbf{x}}_{t}$. For example, for classification tasks, $M[\cdot,\cdot]$ denotes error rate. For question answering tasks, $M[\cdot,\cdot]$ represents the negative value of EM score.
We apply the importance sampling trick to connect the expected $\operatorname{Error}$ and decompose the distributional differences into two-component weights as follows:
$\mathbb{E}_{P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right],$
$=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right]\frac{P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})},$
$=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right](\bm{w}+\bm{\beta}).$
where $P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})/P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})$ measures the distributional difference, $\bm{w}=P_{t}(Y)/P_{c}(Y)$ denotes class-wise weights and $\bm{\beta}=\frac{P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{y}})}\left(\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1\right)$ is conditional bias. The detailed proof can be found in Appendix A.2.
Notably, these existing rebalance methods (i.e. re-weighting (Cui et al., 2019)) assume a shared conditional distribution between annotated and test datasets $P_{c}(X|Y)=P_{t}(X|Y)$ and only focus on how to use the class-wise weights $\bm{w}$ to process imbalanced datasets. However, in the real world, the conditional distribution $P(X|Y)$ is also prone to changes from the annotated dataset to the test dataset due to target shift and concept shift, especially in imbalanced datasets (Gopalan et al., 2013; Jamal et al., 2020; Gu et al., 2022).
We also conduct experiments to verify whether there is a shared conditional distribution between annotated and test datasets $P_{c}(X|Y)=P_{t}(X|Y)$ in imbalanced datasets. Specifically, for each class, we firstly transform text input from each class into bag-of-words representation, employ the Latent Dirichlet Allocation (LDA) model to derive the document-topic distributions and then compute Kullback-Leibler (KL) divergence between annotated conditional distribution $P_{c}(X|Y)$ and test conditional distribution $P_{t}(X|Y)$. For example, as shown in Figure LABEL:figure:4, the KL divergence between annotated and test datasets is unequal to zero and the assumption that $\beta=0$ is difficult to be satisfied.
In this paper, we relax the above assumption and allow a different conditional distribution of annotated and test datasets (e.g., $P_{c}(X|Y)\neq P_{t}(X|Y)$). However, directly computing the conditional bias $\bm{\beta}$ is non-trivial as real conditional distributions of annotated and test datasets are unknown. To circumvent the issue, we aim to design an effective method to estimate the conditional bias $\bm{\beta}$.
4.2 Modeling the conditional bias $\bm{\beta}$
Motivated by the previous analysis, we employ two-component weights, including class-wise weights $\bm{w}$ and conditional bias $\bm{\beta}$, to modify the original scoring functions during selection. We let the class-wise weights resemble the empirically successful design in the literature. Our key idea is to estimate conditional bias $\bm{\beta}$ by achieving empirical $\operatorname{Error}$ minimization of a balanced validation set.
Constructing balanced validation set
Given an imbalanced dataset $\mathcal{D}_{c}$ with $N$ examples, the balanced validation set $\mathcal{D}_{b}$ is constructed by randomly selecting $n_{b}$ examples from $j$-class dataset $\mathcal{D}_{j}$ as follows:
$\mathcal{D}_{b}=\cup_{j=1}^{k}\left\{\operatorname{RandomSample}(\mathcal{D}_{j},n_{b})\right\},$ (7)
where $\operatorname{RandomSample}(\mathcal{D}_{j},n_{b})$ denotes randomly selecting $n_{b}$ examples from each class set $\mathcal{D}_{j}$. Notably, $n_{b}$ is constrained by the condition $n_{b}<\min(n_{1},...,n_{k})$. This process results in a validation set that is balanced across all classes, with each class contributing the same number of examples, similar to the composition of the test dataset. Meanwhile, $\mathcal{D}_{r}$ denotes the imbalanced remaining dataset.
[…]
$K$ 8 16
Imbalance Ratio 0 10 50 100 0 10 50 100
Random 32.63$\pm$0.73 32.26$\pm$1.32 30.91$\pm$1.20 30.21$\pm$0.83 35.25$\pm$0.46 36.17$\pm$0.91 33.24$\pm$1.24 32.51$\pm$0.65
+Ours 32.44$\pm$0.92 36.66$\pm$1.05 36.45$\pm$1.26 37.63$\pm$1.79 34.68$\pm$0.53 36.12$\pm$0.99 38.74$\pm$1.36 40.32$\pm$1.79
TopK 49.90$\pm$0.05 46.51$\pm$0.76 41.51$\pm$1.11 39.62$\pm$1.22 52.31$\pm$0.09 48.18$\pm$0.98 43.34$\pm$0.91 40.95$\pm$0.61
+Ours 49.75$\pm$0.41 49.62$\pm$0.62 46.48$\pm$1.07 44.44$\pm$1.40 51.73$\pm$0.49 51.19$\pm$1.08 48.21$\pm$1.01 45.52$\pm$1.40
DPP 50.56$\pm$0.06 46.59$\pm$0.79 42.54$\pm$1.09 40.26$\pm$1.22 53.66$\pm$0.15 48.38$\pm$0.98 43.53$\pm$0.91 41.83$\pm$0.64
+Ours 50.36$\pm$0.31 49.47$\pm$1.57 46.50$\pm$1.26 45.16$\pm$1.64 52.84$\pm$0.27 50.97$\pm$1.14 48.16$\pm$1.18 46.77$\pm$1.48
VoteK 48.69$\pm$0.14 45.71$\pm$0.88 41.51$\pm$1.13 39.94$\pm$0.95 52.62$\pm$0.24 48.02$\pm$0.71 42.72$\pm$0.91 40.57$\pm$0.60
+Ours 48.57$\pm$0.39 48.21$\pm$0.97 46.39$\pm$1.23 44.77$\pm$1.27 52.62$\pm$0.44 50.97$\pm$0.98 48.05$\pm$0.82 45.84$\pm$1.05
ConE 48.08$\pm$0.86 44.55$\pm$1.51 40.35$\pm$1.29 38.03$\pm$1.31 51.18$\pm$0.58 46.09$\pm$0.79 41.30$\pm$0.63 38.92$\pm$0.72
+Ours 48.51$\pm$0.70 46.89$\pm$0.85 45.09$\pm$1.08 43.21$\pm$1.32 50.95$\pm$0.46 48.32$\pm$0.84 46.83$\pm$1.07 44.13$\pm$0.79
ByDC 50.07$\pm$0.45 45.00$\pm$0.78 41.10$\pm$1.23 38.90$\pm$1.36 53.15$\pm$0.44 46.16$\pm$0.75 41.90$\pm$0.76 38.92$\pm$0.62
+Ours 49.82$\pm$0.39 48.54$\pm$0.94 46.60$\pm$1.11 44.52$\pm$1.05 52.61$\pm$0.55 49.98$\pm$1.00 47.33$\pm$1.23 44.39$\pm$1.14
Bayesian optimizing conditional bias
In practice, since the data distribution is unknown, we use empirical $\operatorname{Error}$ minimization of the balanced validation set $\mathcal{D}_{b}$ to achieve an empirical estimate of the underlying data distribution.
Typically, we minimize the empirical $\operatorname{Error}$ as follows:
$\frac{1}{|\mathcal{D}_{b}|}\sum_{i=1}^{|\mathcal{D}_{b}|}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\{(w_{i}+\beta)\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{b})\right\}_{i=1}^{|\mathcal{D}_{r}|}),\mathrm{\mathbf{x}}_{b}\right),\mathrm{\mathbf{y}}_{b}\right].$
where we still use effective numbers (Cui et al., 2019) as class-wise weights where $w_{i}=(1-\alpha)/(1-\alpha^{n_{j}})$ and $\alpha=(N-1)/N$. We estimate the conditional bias $\beta$ by employing a Bayesian optimizing framework, which is suitable for non-differentiable and black-box function optimization. In this way, the estimated conditional bias $\beta$ can be regarded as an approximation of the conditional bias between annotated and test datasets. We present our overall optimizing process in Appendix A.3.
Re-weighting with adjusted class-wise weights
We use existing selection methods to select a subset of candidates (e.g. $K^{{}^{\prime}}$=100). We sort all candidates in increasing order by $(w_{i}+\beta)\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})$ and obtain the indices for the sorted scores as $\mathcal{I}$. And we select these candidates with low-ranking examples in the sorted list $\mathcal{I}$:
$g(\mathrm{\mathbf{c}}_{i})=\mathrm{1}\left(\operatorname{Loc}(\mathrm{\mathbf{c}}_{i},\mathcal{I})\leq K\right),$ (8)
$\mathcal{I}=\operatorname{argsort}\left\{(w_{i}+\beta)s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t}))\right\}_{i=1}^{K^{{}^{\prime}}}.$ (9)
[…]
  • Flexible: Our method can be easily incorporated into existing selection methods under the class-imbalanced setting in $\mathcal{D}_{c}$.
  • Easy to use: Our method does not require heavy hyperparameter tuning, as it is insensitive to balance set $\mathcal{D}_{b}$ or class-wise weights $\bm{w}$ (see Figure LABEL:figure:5 (a) and (b)).
  • Model-agnostic: Our method requires access only to the model outputs and integrates effortlessly with any LLMs (see Figure LABEL:figure:5 (c) and (d)).
5 Experiments
5.1 Experimental Setup
Datasets and Evaluation For our evaluations, we verify the effectiveness of our method on four benchmark datasets, including Sentiment Classification (Amazon (Keung et al., 2020), Yelp (Zhang et al., 2015)) and Topic Classification (AgNews, Yahoo (Zhang et al., 2015)). Due to limited space, these tasks’ input/output, statistics and split are reported in the Appendix A.4. To simulate the issue of imbalance data, we generate imbalanced datasetss with pre-defined imbalance ratios $\phi$ (e.g., 0, 10, 50, 100) and we evaluate the performance of ICL on another corresponding class-balanced test dataset. We report average accuracy (%) with standard deviation (over 3 runs).
Models and Baselines We utilize various large language models, including Llama3-8B (AI@Meta, 2024), Mistral-7B (Jiang et al., 2023) and OPT-6.7B (Zhang et al., 2022a). We use Bert-base-uncased sentence encoder (Devlin et al., 2019) as the similarity tokenizer. We also conduct experiments with existing selection methods as baselines, including Random (Min et al., 2022), TopK (Liu et al., 2022), DPP (Ye et al., 2023a), VoteK (Su et al., 2023), ConE (Peng et al., 2024), ByCS (Wang et al., 2024). For hyperparameters, we set the size of the balanced validation set as $|\mathcal{D}_{b}|=100$ by default. The details of our experiments are presented in Appendix A.4.
5.2 Main Results
Can our method improve the performance of existing selection methods? Table 1 presents the average test accuracy of in-context learning with different selection methods on four benchmark datasets using OPT-6.7B, under various imbalance ratios. A salient observation is that our method drastically improves the performance of ICL with imbalanced annotated datasets. For example, for the 100 imbalance ratio, our method improves the average accuracy of six existing selection methods from 37.93 to 43.29 – a 5.46 of direct improvement. More importantly, we show that our method can boost performance for a wide range of selection methods, such as TopK and DPP. For example, we observe that, for the 100 imbalance ratio, the test accuracy of the TopK is improved to 4.57 when integrating our methods. Table 1 also presents the results for different demonstration numbers $K$, suggesting that our method can consistently improve the performance of ICL against imbalanced datasets with varying demonstration numbers $K$. The detailed results are reported in Appendix A.6.3.
How does the balanced validation set affect the robustness of our method? In Figure LABEL:figure:5 (a), we ablate how the size of balanced validation set $\mathcal{D}_{b}$ affects the effectiveness of our method (cf. Eq. 7). The baselines indicates all candidate demonstrations are selected without our method. It’s noteworthy that our method shows robustness to the choice of the size of the balanced validation set $|\mathcal{D}_{b}|$. Even when we set $|\mathcal{D}_{b}|=50$, it still yields significant improvements in ICL performance on four classification tasks across six selection methods against imbalanced datasets. Due to space constraints, we only report the average results of multiple baselines on various downstream tasks.
Is our method effective with different class-wise weights $\bm{w}$? In the paper, we employ effective numbers as class-wise weights $\bm{w}$, which have been confirmed to empirically successfully deal with imbalance problems in previous studies (Cui et al., 2019; Jamal et al., 2020). We also verify whether the type of class-wise weights affects the performance of our methods. Specifically, we employ class frequency $w_{j}=\frac{n_{j}}{N}$ (Kang et al., 2020; Shi et al., 2023) as an alternative class-wise weights to verify the effectiveness of our method. Our results in Figure LABEL:figure:5 (b) show that our method can improve the ICL’s performance across various imbalanced ratios and achieve similar accuracy to the effective numbers.
Is our method effective with different model architectures? To demonstrate that our proposed method is model-agnostic, we conduct experiments on diverse open-sourced LLMs (Llama3-8B (AI@Meta, 2024) and Mistral-7B (Jiang et al., 2023)) and present the average results of four common benchmarks using the two LLMs in Figure LABEL:figure:5 (c). From the results, we observe that our method consistently improves test performance on various downstream tasks when using different model architectures. For instance, using our method boosts the test accuracy from 47.10 to 54.31, a 7.21 of direct improvement with a 100 imbalance ratio. Therefore, our method can improve the performance of ICL using various open-source LLMs against imbalanced annotated datasets. More details are presented in Appendix A.6.4.
Is our method effective with closed-sourced LLMs? While our method has demonstrated strong promise in open-source LLMs, we also verify its effectiveness using closed-source models. Specifically, we tested our method on ChatGPT-3.5-Turbo across four selection methods (Random, TopK, DPP and VoteK) with various imbalance ratios. Figure LABEL:figure:5 (d) reports the average results for ChatGPT-3.5-Turbo, showing that imbalanced annotated datasets significantly hurt the ICL performance of closed-source LLMs when employing these existing selection methods. After integrating our method, Figure LABEL:figure:5 (d) also shows significant improvements in average test accuracy across four downstream tasks, indicating the effectiveness of our method in addressing imbalances in closed-source LLMs.
Can our method improve real-world imbalanced dataset We verify the effectiveness of our method on a real-world imbalanced dataset. The Emotion (Saravia et al., 2018) dataset has a long-tailed distribution (refer to Figure 1). Figure 18 shown that our method consistently improves existing selection methods and outperforms existing rebalance methods, including over-sampling (Chawla et al., 2002), under-sampling (Liu et al., 2008), stratified sampling (Vilariño et al., 2005), and re-weighting (Cui et al., 2019). For example, with OPT-6.7B (Zhang et al., 2022a), using our method boosts the average accuracy of six selection methods from 21.22 to 23.36, a relative improvement of 10.08%. These results verify that our method is effective in improving ICL’s performance in real-world imbalanced scenarios.
6 Discussion
6.1 Data imbalance in text generation tasks
[…]
Table 2 demonstrates that imbalanced annotated dataset significantly hurt ICL’s performance on generation tasks. Additionally, Table 2 also confirms the ineffectiveness of classical data balancing methods, including over-sampling (Chawla et al., 2002), under-sampling (Liu et al., 2008), stratified sampling (Vilariño et al., 2005), and re-weighting (Cui et al., 2019). After integrating our method, the existing selection methods show significant improvements in inference performance, indicating the generalization of our method to text generation. For example, our method improves the average exact match score from 21.20 to 22.93, a notable improvement of 0.73 over six existing demonstration methods. The detailed results of CodeSearchNet (Husain et al., 2019) task are presented in Table 7.
Table 2: The average EM score of NQ dataset across six selection methods with various imbalance ratios. Bold numbers are the best results.
Imbalance Ratio 0 10 50 100
Baselines 23.67 22.47 21.33 21.20
+Over-Sampling 23.60 22.53 22.53 22.33
+Under-Sampling 23.60 19.87 18.47 17.80
+Stratified-Sampling 22.15 20.60 18.82 18.47
+Re-Weighting 23.33 22.60 20.47 20.20
+Ours 24.00 23.87 23.00 22.93
6.2 Does a larger dataset address the imbalance issue?
Table 3: Average classification accuracy (%) of Baselines / +Ours methods on the Yahoo dataset with various imbalance ratios. Bold numbers are superior results.
Imbalance Ratio 10 50 100
Baseline/+Ours
Random 37.87/38.07 36.20/39.87 33.13/37.73
TopK 47.33/48.87 45.00/47.67 43.07/47.87
DPP 49.60/50.27 48.00/49.33 44.47/47.80
VoteK 47.13/47.80 43.87/47.47 44.67/47.73
ConE 47.53/49.13 45.27/46.80 41.47/45.53
ByDC 48.40/49.33 44.27/47.47 42.17/46.50
Average 46.64/47.25 43.77/46.44 41.50/45.53
While our analysis has shown that imbalanced annotated datasets negatively impact in-context learning by reducing the number of examples in tail classes, one might wonder if this negative effect is due to the reduced size of the dataset rather than a shift in class prior distribution. To address this, we verify the negative impact of imbalanced annotated datasets and the effectiveness of the proposed method on such datasets with an increasing number of examples. Specifically, we consider creating an imbalanced annotated dataset by increasing the number of examples in head classes while keeping the number of examples in tail classes constant. For example, in the case of Yahoo (Zhang et al., 2015), when the imbalance ratio $\phi=0$, the number of annotated examples for each class remains 50. When the imbalance ratio $\phi=100$, the number of examples in the head classes is 5,000, but the number in the tail classes remains 50. We report the test accuracy of Yahoo using OPT-6.7B with six existing selection methods.
[…]
7 Conclusion
In this paper, we study the effect of the imbalanced annotated dataset on the performance of in-context learning (ICL). Both our theoretical analysis and empirical investigations reveal that imbalanced annotated datasets significantly hurt the performance of ICL. Meanwhile classical rebalance methods cannot effectively mitigate the imbalanced datasets. To address the issue, we introduce our method, a general strategy that can universally enhance the performance of ICL across various imbalanced ratios. Our method is motivated by decomposing the distributional differences between annotated and test datasets into class-wise weights and conditional bias. Extensive experiments demonstrate that our method can improve the performance of existing selection methods of ICL across various imbalance ratios. Our approach is practical to use, as it is insensitive to the hyperparameters and integrates effortlessly with any LLMs.
Limitations. Our methods need to split a balanced validation dataset from imbalanced annotated datasets, and these methods require that the size of examples in tail classes cannot be too small. Furthermore, it might be challenging to estimate multiple parameters for the class-conditional biases using balanced datasets with limited size. It might be an interesting direction to explore how to model the bias in a more data-efficient manner.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 Related Work
In-context learning In-context learning (ICL) is a new paradigm for large language (LLMs), which allows LLMs to make predictions only based on a few demonstrations without explicitly updating parameters (Akyurek et al., 2023; Hendel et al., 2023; Agarwal et al., 2024; Dong et al., 2024; Edwards & Camacho-Collados, 2024; Falck et al., 2024). Many studies show that ICL can achieve performance similar to fine-tuning but without the high computational cost (Gonen et al., 2023; Mosbach et al., 2023; Müller et al., 2024; Panwar et al., 2024). Despite achieving such outstanding performance, ICL has been criticized for being very sensitive to the quality of in-context examples (Fei et al., 2023; Gao et al., 2024a). Various approaches have been proposed to improve the robustness of ICL in recent years, including meta-tuning LLMs (Brunet et al., 2023), calibration (Abbas et al., 2024), demonstration selection (Zhang et al., 2022b; Nguyen & Wong, 2023; Qin et al., 2024; Ye et al., 2023b; Gao et al., 2024b; Luo et al., 2024; Mo et al., 2024), ordering (Lu et al., 2022; Liu et al., 2024), number (Zhang et al., 2025) and formation (Voronov et al., 2024; Yao et al., 2024).
Notably, existing studies often assume that usually assume that annotated dataset and testing data are both i.i.d. (independent and identically distributed) sampled from the same data distribution (Luo et al., 2024; Van et al., 2024). However, this is not always true for real-world scenarios. Specifically, annotated dataset typically exhibit a long-tailed class distribution, where a small portion of classes have a massive number of sample points but the others are associated with only a few examples (Jamal et al., 2020; Schultheis et al., 2024). It is still mysterious how demonstrations sampled from long-tailed distribution affect the performance of ICL on both text classification and generation tasks. In this paper, we show that annotation quality is crucial for ICL in both text classification and generation, where imbalanced datasetss significantly hurt the performance. Different from previous study (Hong et al., 2024), we propose a simple and effective method to enhance the performance of ICL from the perspective of demonstration selection.
Learning with imbalanced datasets Imbalanced dataset is common in many real-world datasets (Cui et al., 2018, 2019; Jamal et al., 2020; Schultheis et al., 2024). The existing approaches to learning with imbalanced datasets can be classified into two types: (1) training imbalance-robust models with imbalance training datasets: designing imbalance-robust loss function (Jamal et al., 2020; Tan et al., 2020; Park et al., 2023; Bhat et al., 2023; Garcin et al., 2022) or designing imbalance-robust model architectures (Long et al., 2022; Pan et al., 2024) to mitigate the issue of imbalanced datasets. However, this method is not suitable for ICL, which usually hypothesizes that users are unable to apply fine-tuning techniques. (2) handling imbalance examples: handling imbalanced datasets is crucial for ensuring balanced performance across all classes. One simple and intuitive approach to deal with the class-imbalanced problem is re-sampling. Under-sampling methods remove examples from the majority classes, which is infeasible under data imbalanced settings (Liu et al., 2008; Wei et al., 2022). The over-sampling method adds repeated examples for the minority classes, usually causing over-fitting to the minority classes (Chawla et al., 2002; Shi et al., 2023). It is noteworthy that these existing methods mainly focus on the fine-tuning setting, and the literature on how to mitigate the effects of imbalanced datasets in in-context learning is limited. This motivated us to design a method for addressing the issue of imbalanced datasets under the setting of ICL.
A.2 Proof
Proof. We apply the importance sampling trick to connect the expected $\operatorname{Error}$ with the imbalanced annotated dataset:
$\operatorname{Error}$ $=\mathbb{E}_{P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right],$
$=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right]\frac{P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})},$
$=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right]\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{c}(\mathrm{\mathbf{y}})},$
Here, we decompose the term $\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{c}(\mathrm{\mathbf{y}})}$ into two parts: the class-wise weights $\bm{w}=$, and the conditional weight $\bm{\widetilde{\beta}}=$. The class-wise weights measure the difference between class prior distributions $P(Y)$, and the conditional weight measures the difference between conditional distributions $P(X|Y)$. Then we have
$\operatorname{Error}=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right]\bm{w}(1+\bm{\widetilde{\beta}})$
where $\bm{w}$ denotes class-wise weights and conditional weight $\bm{\widetilde{\beta}}=\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1$ measures the difference of conditional distribution between the annotated and test datasets.
For simplicity, we introduce a conditional bias $\bm{\beta}=\bm{w}\times\bm{\widetilde{\beta}}$ and re-write the expected $\operatorname{Error}$ as:
$\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right](\bm{w}+\bm{\beta}).$ (10)
where conditional bias $\bm{\beta}=\bm{w}\times\bm{\widetilde{\beta}}=\frac{P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{y}})}\left(\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1\right)$ denotes the deviation of class-wise weights caused by the presence of conditional weight $\bm{\widetilde{\beta}}$.
A.3 Bayesian Optimizing Conditional Bias
In this section, we will introduce how to use Bayesian optimization (Gardner et al., 2014; Nogueira, 2014) to find the optimal value of conditional bias $\beta$ by maximizing empirical $\operatorname{Error}$ as follows:
$\mathcal{F}_{\theta}\left(\beta|\mathcal{D}_{r},\mathcal{D}_{b},\bm{w}\right)=\frac{1}{|\mathcal{D}_{b}|}\sum_{i=1}^{|\mathcal{D}_{b}|}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{(w_{i}+\beta)\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{b})\right\}_{i=1}^{|\mathcal{D}_{r}|}\right),\mathrm{\mathbf{x}}_{b}\right),\mathrm{\mathbf{y}}_{b}\right].$ (11)
Surrogate Model. We use the Gaussian Process as a surrogate model to approximate the objective function 11. Initially, a prior distribution is established in a general form for the estimated model parameters,
$\mathcal{F}_{\theta}\left(\beta|\mathcal{D}_{r},\mathcal{D}_{b},\bm{w}\right)=\mathcal{N}\left(\mu(\beta),\sigma^{2}(\beta)\right),$
where $\mathcal{N}$ is the Gaussian distribution with a mean function $\mu(\beta)$ and a covariance function $\sigma^{2}(\beta)$. For a data point $\beta_{j}$ sampled from the Gaussian process $\mathcal{N}$, we compute its corresponding function values $\mathcal{F}_{\theta}\left(\beta_{j}|\mathcal{D}_{r},\mathcal{D}_{b},w\right)$.
Acquisition Function. We use the Expected Improvement (EI) criterion to select the point $\beta_{j+1}$ with the maximum expected improvement as the next query point:
$\operatorname{EI}(\beta)$ =$=\left\{\begin{array}[]{ll}\left(\mu_{j}(\beta)-\mathcal{F}_{\theta}\left(\beta^{+}|\mathcal{D}_{r},\mathcal{D}_{b},w\right)-\epsilon\right)\Phi(Z)+\sigma_{j}(\beta)\phi(Z)&\text{ if }\sigma_{j}(\beta)>0,\\ 0&\text{ if }\sigma_{j}(\beta)=0.\end{array}\right.$ (13)
$Z$ $=\frac{\mu_{j}(\beta)-\mathcal{F}_{\theta}\left(\beta^{+}|\mathcal{D}_{r},\mathcal{D}_{b},w\right)-\epsilon}{\sigma_{j}(\beta)}.$
where $\mu_{j}(\beta)$ and $\sigma_{j}(\beta)$ represent the mean and standard deviation of the surrogate model at step $j$, $\beta^{+}$ denotes the current optimal observed point, $\epsilon$ is a tunable hyper-parameter. $\Phi(Z)$ and $\phi(Z)$ are the probability density function and cumulative density function of the Gaussian Process, respectively.
Iterative Optimization Process. We start with an initial set of observations $\beta_{0},...,\beta_{h}$ by evaluating the objective function $\left\{\mathcal{F}\left(\beta_{i}|\mathcal{D}_{r},\mathcal{D}_{b},w,\theta\right)\right\}_{i=0}^{h}$ at a few selected points. Second, we use the observed data $\left(\beta_{j},\mathcal{F}\left(\beta_{j}|\mathcal{D}_{r},\mathcal{D}_{b},w,\theta\right)\right)$ to update the Gaussian Process model, refining the estimates of $\mu_{j}(\beta)$ and $\sigma_{j}(\beta)$. Third, we employ the Expected Improvement criterion to determine the next point $\beta_{j+1}$ and obtain $\left\{\mathcal{F}\left(\beta_{j+1}|\mathcal{D}_{r},\mathcal{D}_{b},w,\theta\right)\right\}_{i=0}^{h}$. We continue the process of updating the model and selecting new points until a stopping criterion is met. Finally, we select the point $\beta^{+}$ with the best observed value of the objective function $\mathcal{F}_{\theta}\left(\beta^{+}|\mathcal{D}_{r},\mathcal{D}_{b},w\right)$ as the optimal solution.
A.4 Experimental Setting
Datasets We conduct experiments on various classification and generation tasks and examples of each dataset are shown in Tables 9 and 10. We collect all datasets from Huggingface. The train sets of these datasets are regarded as example datasets and the test datasets are used to evaluate the performance of ICL. We randomly subsample examples from the test dataset to verify the performance of the imbalanced annotated dataset. We generate categories for Natural question (Kwiatkowski et al., 2019) using ChatGPT-3.5-Turbo. The categories include people, time, geography, culture, and specialized knowledge.
Baselines To verify the effectiveness of our method, we compare our method with previous methods for demonstration retrieval by the downstream ICL’s performance. We consider both learning-free and other learning-based retrievers as baselines, including
  1. 1.
    Random selects demonstrations randomly from an example set without repetition (Min et al., 2022).
  2. 2.
    TopK retrieves demonstrations that are semantically similar to a test query sample (Liu et al., 2022).
  3. 3.
    DPP uses the original BERT embeddings as mentioned above without fine-tuning, and adopts MAP inference for subset retrieval (Ye et al., 2023a).
  4. 4.
    VoteK proposes an unsupervised and graph-based selective annotation method to select diverse and representative demonstrations (Su et al., 2023).
  5. 5.
    ConE searches for demonstrations by minimizing the difference in cross-entropy between the test input and the demonstrations (Peng et al., 2024).
  6. 6.
    ByCS assumes that an accurate inverse likelihood probability will lead to an accurate posterior probability and selects demonstrations based on their inverse inference results (Wang et al., 2024).
Experiment details We run our experiments on NVIDIA GeForce RTX 4090 and NVIDIA L40 GPU, and implement all methods by PyTorch and transformers. Our code is inspired by OpenICL (Wu et al., 2023). We thank the authors for releasing their code.
A.5 Inference
Perplexity For classification tasks, we compute the sentence perplexity for each sequence formed by concatenating the input with each candidate answer (Brown et al., 2020; Wu et al., 2023). Specifically, for each input instance $\mathrm{\mathbf{x}}$ is paired potential label set $\mathcal{Y}$, where $\mathcal{Y}$ represents the set of possible classes (e.g., Sports, Business, etc.). Then, for each possible label $\mathrm{\mathbf{y}}\in\mathcal{Y}$, we concatenate each tokenized input-output pair $(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})$, and obtain the corresponding tokenized sequence $\mathrm{\mathbf{c}}=(z_{1},...,z_{|\mathrm{\mathbf{c}}|})=(x_{1},...,x_{|\mathrm{\mathbf{x}}|},y_{1},...,y_{|\mathrm{\mathbf{y}}|})$, where $|\mathrm{\mathbf{c}}|=|\mathrm{\mathbf{x}}|+|\mathrm{\mathbf{y}}|$. Now, the perplexity of $\mathrm{\mathbf{c}}$ is calculated as:
$\operatorname{Perplexity}(\mathrm{\mathbf{c}})=\operatorname{exp}\{-\frac{1}{|\mathrm{\mathbf{c}}|}\sum^{|\mathrm{\mathbf{c}}|}_{i=1}\log p_{\theta}(c_{i}|c_{<i})\}.$
where $\log p_{\theta}(c_{i}|c_{<i})$ is the log-likelihood of the $i$-th token conditioned on the preceding tokens $c_{<i}$, from the given language model parameterized by $\theta$. We select the label corresponding to the input-output pair $\mathrm{\mathbf{c}}$ with the lowest perplexity as the predicted label for the input $\mathrm{\mathbf{x}}$.
Direct For text generation tasks, we represent candidate answers using tokens from the vocabulary and select the final prediction based on the one with the highest probability (Brown et al., 2020; Wu et al., 2023). To reduce computational cost, we set the maximum new tokens to 50 and limit unnecessary token generation.
A.6 More empirical results
A.6.1 The effect of imbalanced dataset
In this section, we show the effect of imbalanced datasets on the performance of in-context learning with different model architectures and various numbers of subsets. Figures LABEL:figure:a1 (b) and (c) report that imbalanced datasets significant degrade ICL’s performance of Llama3-8B (AI@Meta, 2024) and Mistral-7B (Jiang et al., 2023). This illustrates that the negative effect of imbalanced datasets can be commonly observed in various large language models. Additionally, Figure LABEL:figure:a1 (a) also reports a larger set of demonstrations does not enhance the performance of ICL. Specifically, imbalanced datasets also negatively impact the performance of in-context learning.
A.6.2 The failure of classical rebalance methods
Figures LABEL:figure:a3 (a) and (b) reports the Kullback-Leibler (KL) divergence between $P_{c}(X|Y)$ and $P_{t}(X|Y)$ for each class in the Yelp and Yahoo datasets, respectively. Notably, the KL divergence between $P_{c}(X|Y)$ and $P_{t}(X|Y)$ for each class is unequal to zero and the assumption that $\beta$ = 0 is difficult to be satisfied.
A.6.3 Main results for each dataset
[…]
A.6.4 Is our method effective with different model architectures?
In this section, we provide the detailed results of Llama3-8B (AI@Meta, 2024) and Mistral-7B (Jiang et al., 2023). Figures LABEL:figure:a4 (a) and (b) report the average accuracy of four classification tasks across six selection methods. We can clearly observe that imbalanced annotated datasets significantly hurt the performance of in-context learning, which is consistent with our previous findings. Meanwhile, Figures LABEL:figure:a4 (a) and (b) also demonstrate that integrating our method can improve the performance of ICL against imbalanced annotated datasets.
A.6.5 Can our method improve real-world imbalanced dataset?
We further verify the effectiveness of our method on a real-world class-imbalanced dataset (Emotion (Saravia et al., 2018)). .We select demonstrations from the training set and construct a balanced test set from the original test set. Figure 18 summarizes the test accuracy across six existing selection methods using baselines, classical rebalance methods, and our method on the Emotion dataset. In particular, the proposed method is able to consistently improve the existing selection methods and outperforms existing rebalance methods in test accuracy.
[…]
$K$ 8 16
Imbalance Ratio 0 10 50 100 0 10 50 100
Random 68.17$\pm$0.12 58.58$\pm$1.25 48.75$\pm$1.81 46.25$\pm$0.35 70.58$\pm$0.59 61.33$\pm$0.82 50.75$\pm$1.62 48.25$\pm$1.24
+Ours 68.08$\pm$0.72 67.75$\pm$1.43 69.75$\pm$0.82 69.92$\pm$1.78 68.83$\pm$0.42 69.83$\pm$0.31 70.01$\pm$2.36 73.42$\pm$4.19
TopK 86.67$\pm$0.12 76.42$\pm$0.77 70.25$\pm$1.27 66.67$\pm$1.89 88.75$\pm$0.00 79.17$\pm$0.42 71.50$\pm$1.08 68.00$\pm$0.35
+Ours 85.92$\pm$1.18 86.17$\pm$0.62 81.83$\pm$1.53 76.50$\pm$1.24 88.33$\pm$0.24 86.58$\pm$1.23 83.25$\pm$1.27 76.42$\pm$1.01
DPP 86.50$\pm$0.00 76.42$\pm$1.83 69.67$\pm$0.82 66.25$\pm$1.34 87.25$\pm$0.00 78.42$\pm$1.36 71.58$\pm$1.71 68.92$\pm$0.31
+Ours 86.42$\pm$0.12 87.75$\pm$1.43 81.25$\pm$1.67 78.17$\pm$1.33 87.42$\pm$0.12 87.42$\pm$1.84 82.83$\pm$1.53 77.50$\pm$2.30
VoteK 85.17$\pm$0.12 76.75$\pm$1.54 69.58$\pm$0.82 66.42$\pm$1.66 86.42$\pm$0.12 78.83$\pm$0.42 72.17$\pm$0.92 68.67$\pm$1.23
+Ours 85.50$\pm$0.20 85.92$\pm$0.24 82.17$\pm$1.71 77.42$\pm$1.50 86.58$\pm$0.31 86.25$\pm$1.41 82.50$\pm$0.00 76.75$\pm$1.41
ConE 85.58$\pm$1.12 76.00$\pm$2.16 69.00$\pm$1.97 65.50$\pm$1.27 85.75$\pm$1.14 77.33$\pm$1.25 71.08$\pm$0.82 66.5$\pm$0.54
+Ours 85.83$\pm$0.94 82.83$\pm$1.31 79.50$\pm$0.54 74.85$\pm$2.46 86.08$\pm$0.72 83.75$\pm$1.08 81.58$\pm$0.62 74.83$\pm$1.36
ByDC 85.40$\pm$0.43 74.00$\pm$1.08 67.58$\pm$1.16 63.58$\pm$1.36 87.92$\pm$0.47 76.08$\pm$0.59 69.58$\pm$1.12 63.75$\pm$0.54
+Ours 84.73$\pm$0.38 82.75$\pm$1.24 80.92$\pm$1.36 77.17$\pm$1.85 88.33$\pm$0.12 86.25$\pm$1.41 81.50$\pm$1.41 72.75$\pm$1.41
Baselines 82.92$\pm$0.32 73.03$\pm$1.44 65.81$\pm$1.31 62.45$\pm$1.31 84.45$\pm$0.39 75.19$\pm$0.81 67.78$\pm$1.21 64.02$\pm$0.70
+Ours 82.75$\pm$0.59 82.20$\pm$1.05 79.24$\pm$1.27 75.67$\pm$1.69 84.26$\pm$0.32 83.35$\pm$1.21 80.28$\pm$1.20 75.28$\pm$1.95
[…]
$K$ 8 16
Imbalance Ratio 0 10 50 100 0 10 50 100
Random 37.93$\pm$1.91 39.53$\pm$2.22 36.07$\pm$0.41 33.13$\pm$1.47 44.20$\pm$0.59 44.47$\pm$0.81 39.27$\pm$1.18 37.60$\pm$0.59
+Ours 37.87$\pm$1.84 38.87$\pm$0.81 36.00$\pm$0.71 37.73$\pm$1.64 43.20$\pm$1.18 42.93$\pm$0.25 44.33$\pm$0.74 44.00$\pm$1.93
TopK 52.80$\pm$0.00 51.47$\pm$0.34 45.53$\pm$0.68 43.07$\pm$1.46 53.67$\pm$0.19 52.93$\pm$1.27 50.27$\pm$0.90 47.27$\pm$1.32
+Ours 53.47$\pm$0.47 53.00$\pm$0.33 48.33$\pm$1.54 47.87$\pm$2.62 54.40$\pm$0.85 56.00$\pm$1.50 50.53$\pm$1.33 50.20$\pm$2.70
DPP 54.00$\pm$0.00 52.80$\pm$0.16 47.53$\pm$1.24 44.47$\pm$2.00 56.57$\pm$0.26 54.03$\pm$0.97 50.47$\pm$0.57 48.87$\pm$1.25
+Ours 53.80$\pm$0.28 51.04$\pm$1.49 48.13$\pm$0.74 47.80$\pm$2.27 56.07$\pm$0.19 54.87$\pm$0.68 51.07$\pm$1.46 51.60$\pm$1.82
VoteK 52.13$\pm$0.19 50.80$\pm$0.59 46.67$\pm$1.09 44.67$\pm$1.31 56.37$\pm$0.45 53.10$\pm$0.54 48.70$\pm$1.08 46.20$\pm$0.82
+Ours 51.73$\pm$0.52 50.24$\pm$0.45 48.53$\pm$1.84 47.73$\pm$1.59 56.03$\pm$0.26 55.10$\pm$1.10 52.03$\pm$1.31 51.07$\pm$0.94
ConE 51.53$\pm$0.52 49.20$\pm$0.86 44.60$\pm$1.30 41.47$\pm$1.89 51.90$\pm$0.80 49.37$\pm$0.90 45.33$\pm$0.68 42.77$\pm$1.47
+Ours 51.67$\pm$0.66 50.07$\pm$0.41 46.93$\pm$0.75 45.53$\pm$1.16 52.90$\pm$0.67 49.37$\pm$0.90 49.00$\pm$1.50 47.43$\pm$0.83
ByDC 53.67$\pm$0.74 49.60$\pm$0.85 44.80$\pm$1.70 42.17$\pm$1.68 55.03$\pm$0.58 49.77$\pm$1.11 47.40$\pm$0.86 44.87$\pm$1.25
+Ours 53.40$\pm$0.43 52.33$\pm$0.90 48.47$\pm$1.65 46.50$\pm$0.57 55.37$\pm$0.68 52.07$\pm$0.94 50.40$\pm$0.86 48.87$\pm$0.94
Baselines 50.34$\pm$0.56 48.90$\pm$0.84 44.20$\pm$1.07 41.50$\pm$1.64 52.96$\pm$0.48 50.61$\pm$0.93 46.91$\pm$0.88 44.60$\pm$1.12
+Ours 50.32$\pm$0.70 49.26$\pm$0.73 45.96$\pm$1.36 45.53$\pm$1.64 53.00$\pm$0.64 51.72$\pm$0.90 49.56$\pm$1.20 48.86$\pm$1.53
[…]
$K$ 8 16
Imbalance Ratio 0 10 50 100 0 10 50 100
Random 12.33$\pm$0.62 15.2$\pm$1.45 18.53$\pm$1.32 20.13$\pm$0.98 13.33$\pm$0.19 18.53$\pm$1.05 20.07$\pm$0.09 20.33$\pm$0.34
+Ours 12.53$\pm$0.47 19.87$\pm$0.96 20.93$\pm$2.38 20.47$\pm$1.33 13.47$\pm$0.25 16.13$\pm$2.54 20.20$\pm$1.18 21.07$\pm$0.62
TopK 28.73$\pm$0.09 28.47$\pm$0.74 24.40$\pm$1.80 23.40$\pm$0.86 30.60$\pm$0.00 28.13$\pm$1.15 24.20$\pm$0.65 23.07$\pm$0.66
+Ours 28.80$\pm$0.00 29.47$\pm$0.50 27.87$\pm$0.96 27.20$\pm$0.49 30.40$\pm$0.00 31.00$\pm$1.30 29.13$\pm$0.93 28.33$\pm$0.90
DPP 29.80$\pm$0.16 27.13$\pm$0.68 25.47$\pm$1.27 24.13$\pm$0.5 32.47$\pm$0.09 28.13$\pm$0.47 24.00$\pm$0.43 23.00$\pm$0.75
+Ours 29.60$\pm$0.00 28.67$\pm$1.79 28.67$\pm$0.81 27.60$\pm$0.86 31.87$\pm$0.75 30.87$\pm$1.39 28.93$\pm$1.00 27.93$\pm$1.06
VoteK 28.67$\pm$0.09 25.67$\pm$0.81 22.73$\pm$1.23 22.93$\pm$0.41 31.57$\pm$0.29 27.67$\pm$0.82 22.60$\pm$0.65 21.93$\pm$0.25
+Ours 28.13$\pm$0.09 27.53$\pm$1.72 27.00$\pm$0.75 27.47$\pm$1.32 31.80$\pm$0.85 30.07$\pm$0.34 28.40$\pm$1.40 27.20$\pm$1.02
ConE 26.73$\pm$0.47 25.40$\pm$1.34 22.87$\pm$0.82 21.27$\pm$1.16 30.20$\pm$0.00 27.73$\pm$0.47 22.87$\pm$0.94 22.27$\pm$0.38
+Ours 27.40$\pm$0.82 26.73$\pm$1.27 27.60$\pm$1.30 27.27$\pm$1.16 30.27$\pm$0.34 29.07$\pm$1.20 29.13$\pm$0.93 26.80$\pm$0.49
ByDC 29.47$\pm$0.52 27.13$\pm$0.34 25.20$\pm$1.02 23.53$\pm$1.57 32.13$\pm$0.38 27.93$\pm$0.66 24.13$\pm$0.50 22.27$\pm$0.25
+Ours 29.67$\pm$0.50 28.80$\pm$1.42 29.20$\pm$0.71 27.13$\pm$1.05 31.53$\pm$1.23 30.07$\pm$1.31 28.60$\pm$1.14 27.8$\pm$0.75
Baselines 25.96$\pm$0.33 24.83$\pm$0.89 23.20$\pm$1.24 22.57$\pm$0.91 28.38$\pm$0.16 26.35$\pm$0.77 22.98$\pm$0.54 22.15$\pm$0.44
+Ours 26.02$\pm$0.31 26.85$\pm$1.28 26.88$\pm$1.15 26.19$\pm$1.04 28.26$\pm$0.57 27.87$\pm$1.35 27.40$\pm$1.10 26.52$\pm$0.81
[…]
$K$ 8 16
Imbalance Ratio 0 10 50 100 0 10 50 100
Random 12.07$\pm$0.25 15.73$\pm$0.34 20.27$\pm$1.25 21.33$\pm$0.52 12.87$\pm$0.47 20.33$\pm$0.96 22.87$\pm$2.07 23.87$\pm$0.41
+Ours 11.27$\pm$0.66 20.13$\pm$0.98 19.13$\pm$1.11 22.40$\pm$2.41 13.20$\pm$0.28 15.60$\pm$0.86 20.40$\pm$1.14 22.80$\pm$0.43
TopK 31.40$\pm$0.00 29.67$\pm$1.20 25.87$\pm$0.68 25.33$\pm$0.66 36.20$\pm$0.16 32.47$\pm$1.06 27.40$\pm$0.99 25.47$\pm$0.09
+Ours 30.80$\pm$0.00 29.83$\pm$1.01 27.87$\pm$1.05 26.20$\pm$1.23 33.60$\pm$0.85 31.17$\pm$0.29 29.93$\pm$0.52 27.13$\pm$0.98
DPP 31.93$\pm$0.09 30.00$\pm$0.49 27.47$\pm$1.04 26.20$\pm$1.02 38.33$\pm$0.25 32.93$\pm$1.11 28.07$\pm$0.93 26.53$\pm$0.25
+Ours 31.60$\pm$0.85 30.40$\pm$1.57 27.93$\pm$0.93 27.07$\pm$2.10 36.00$\pm$0.00 30.73$\pm$0.66 29.80$\pm$0.71 30.03$\pm$0.73
VoteK 28.80$\pm$0.16 29.60$\pm$0.59 27.07$\pm$1.39 25.73$\pm$0.41 36.13$\pm$0.09 32.47$\pm$1.06 27.40$\pm$0.99 25.47$\pm$0.09
+Ours 28.93$\pm$0.74 29.13$\pm$1.48 27.87$\pm$0.62 26.47$\pm$0.66 36.07$\pm$0.34 32.47$\pm$1.06 29.27$\pm$0.57 28.33$\pm$0.82
ConE 28.47$\pm$1.32 27.60$\pm$1.66 24.93$\pm$1.05 23.87$\pm$0.93 36.87$\pm$0.38 29.93$\pm$0.52 25.93$\pm$0.09 24.13$\pm$0.47
+Ours 29.13$\pm$0.38 27.93$\pm$0.41 26.33$\pm$1.72 25.20$\pm$0.49 34.53$\pm$0.09 31.07$\pm$0.19 27.60$\pm$1.23 27.47$\pm$0.47
ByDC 31.73$\pm$0.09 29.27$\pm$0.84 26.80$\pm$1.02 26.33$\pm$0.84 37.53$\pm$0.34 30.87$\pm$0.62 26.47$\pm$0.57 24.80$\pm$0.43
+Ours 31.47$\pm$0.25 30.27$\pm$0.19 27.80$\pm$0.72 27.27$\pm$0.74 35.20$\pm$0.16 31.53$\pm$0.34 28.80$\pm$1.50 28.13$\pm$1.46
Baselines 27.40$\pm$0.32 26.98$\pm$0.85 25.40$\pm$1.07 24.80$\pm$0.73 32.99$\pm$0.28 29.83$\pm$0.89 26.36$\pm$0.94 25.05$\pm$0.29
+Ours 27.20$\pm$0.48 27.95$\pm$0.94 26.16$\pm$1.02 25.77$\pm$1.27 31.43$\pm$0.29 28.76$\pm$0.57 27.63$\pm$0.95 27.32$\pm$0.82
Table 8: The average Rouge score of CodeSearchNet dataset across six selection methods with various imbalance ratios. Bold numbers are the best results.
Imbalance Ratio 0 10 50 100
Baselines 15.99 13.26 11.41 9.84
+Over-Sampling 16.12 13.82 10.07 8.56
+Under-Sampling 16.08 10.17 10.06 9.20
+Stratified-Sampling 10.59 9.95 9.54 9.24
+Re-Weighting 15.86 13.33 8.59 7.96
+Ours 15.71 13.64 11.86 10.90
Figure 18: Average accuracy of Emotion tasks across six existing selection methods using OPT-6.7B (Zhang et al., 2022a).
Table 9: The statistics, split and evaluation metrics of each dataset.
Data Train Set test dataset Classes Evaluation
Amazon 25000 1000 5 Accuracy
AgNews 20000 1000 4 Accuracy
Yelp 25000 1000 5 Accuracy
Yahoo 50000 500 10 Accuracy
Emotion 20000 300 6 Accuracy
Natural Question 25000 500 5 Exact Match
CodeSearchNet 18000 500 6 Rouge
Table 10: The instructions, inference templates and example cases of tasks.
Dataset Prompt Example
Amazon Task Instruction: Sentiment of the sentence Inference Verbalizer: Great, Good, Okay, Bad, Terrible? Input: Question Output: Answer Task Instruction: Sentiment of the sentence Inference Verbalizer: Great, Good, Okay, Bad, Terrible? Input: why give me a date for a month then when its suppose to ship, its running late Output: Terrible
AgNews Task Instruction: Text Classification Task Inference Verbalizer: World, Sports, Business or Science New Topic? Input: Question Output: Answer Task Instruction: Text Classification Task Inference Verbalizer: World, Sports, Business or Science New Topic? Input: EBay Buys 25 Percent Stake in Craigslist Network By MAY WONG SAN JOSE, Calif. (AP) – Online auctioneer eBay Inc… Output: Science
Yelp Task Instruction: Sentiment of the sentence Inference Verbalizer: Great, Good, Okay, Bad, Terrible? Input: Question Output: Answer Task Instruction: Sentiment of the sentence Inference Verbalizer: Great, Good, Okay, Bad, Terrible? Input: Awesome place!!! You must go and try all the services!!!! Output: Good
Yahoo Topic of the text: Society &\&& Culture, Science &\&& Mathematics, Health, Education &\&& Reference, Computers &\&& Internet, Sports, Business &\&& Finance, Entertainment &\&& Music, Family &\&& Relationships, Politics &\&& Government? Input: Question Output: Answer Topic of the text: Society &\&& Culture, Science &\&& Mathematics, Health, Education &\&& Reference, Computers &\&& Internet, Sports, Business &\&& Finance, Entertainment &\&& Music, Family &\&& Relationships, Politics &\&& Government? Input: what is god’s kingdom that we are told to pray for? Output: Computers &\&& Internet
Emotion Task Instruction: Sentiment of the sentence Inference Verbalizer: Sadness, Joy, Love, Anger, Fear, Surprise? Input: Question Output: Answer Task Instruction: Sentiment of the sentence Inference Verbalizer: Sadness, Joy, Love, Anger, Fear, Surprise? Input: im feeling generous this week Output: Joy
NQ Question: Question Answer: Answer Question: who is the CEO of what’s up Answer: Jan Koum
[…]

### v2

@@
 Large language models (LLMs) have shown impressive performance on downstream tasks through in-context learning (ICL), which heavily relies on the demonstrations selected from annotated datasets.
-Existing selection methods may hinge on the distribution of annotated datasets, which can often be long-tailed in real-world scenarios.
-In this work, we show that imbalanced class distributions in annotated datasets significantly degrade the performance of ICL across various tasks and selection methods.
-Moreover, traditional rebalance methods fail to ameliorate the issue of class imbalance in ICL.
-Our method is motivated by decomposing the distributional differences between annotated and test datasets into two-component weights: class-wise weights and conditional bias.
-The key idea behind our method is to estimate the conditional bias by minimizing the empirical error on a balanced validation dataset and to employ the two-component weights to modify the original scoring functions during selection.
-Our approach can prevent selecting too many demonstrations from a single class while preserving the effectiveness of the original selection methods.
-Extensive experiments demonstrate the effectiveness of our method, improving the average accuracy by up to 5.46 on common benchmarks with imbalanced datasets.
-Machine Learning, ICML
+However, these datasets often exhibit long-tailed class distributions in real-world scenarios, leading to biased demonstration selection.
+In this work, we show that such class imbalances significantly degrade the ICL performance across various tasks, regardless of selection methods.
+Moreover, classical rebalancing methods, which focus solely on class weights, yield poor performance due to neglecting condition bias–skewed feature distributions within classes.
+To address this, we propose Reweighting with Conditional Bias (dubbed RCB), a simple and complementary approach to enhance ICL performance under class imbalance.
+In particular, RCB estimates conditional bias using a balanced subset and re-weights demonstration scores based on both class weight and conditional bias.
+In effect, RCB prevents over-selection from dominant classes while preserving the efficacy of current selection methods.
+Extensive experiments on common benchmarks demonstrate the effectiveness of our method, improving the average accuracy of current selection methods by up to 5.42%.
 1 Introduction
 In-context learning (ICL) has been a remarkable capability of Large Language Models (LLMs), enabling them to perform downstream tasks by simply conditioning on a few task demonstrations.
-(Brown et al., 2020).
-ICL consistently outperforms zero-shot inference across various tasks without needing parameter updates, positioning it as a strong alternative to supervised fine-tuning (SFT) (Mosbach et al., 2023; Panwar et al., 2024).
-In particular, previous studies typically focus on designing methods for demonstration selection from high-quality annotated datasets (Liu et al., 2022; Baldassini et al., 2024; Wang et al., 2024).
-The effectiveness of these selection algorithms may hinge on the distribution of annotated datasets, which can often be long-tailed in real-world scenarios (Wei et al., 2022; De Alvis & Seneviratne, 2024) — certain classes are underrepresented (see Figure 1).
-Thus, it is of great importance to investigate how such imbalances affect ICL’s performance.
-Figure 1: An example of imbalanced dataset.
-In Emotion (Saravia et al., 2018), a few sentiments make a large contribution and data tend to show a long-tailed distribution.
-For example, “joy“ and “sadness“ are head classes, while most other classes, such as “love“ and “surprise“, are tail classes.
+[1].
+ICL consistently outperforms zero-shot inference across various tasks without needing parameter updates, positioning it as a strong alternative to supervised fine-tuning (SFT) [2, 3].
+Previous studies typically focus on designing methods for demonstration selection from high-quality annotated datasets [4, 5, 6].
+The effectiveness of these selection algorithms may hinge on the distribution of annotated datasets, which can often be long-tailed in real-world scenarios [7, 8] — certain classes are underrepresented (see Figure 5).
+Thus, it is of great importance to explore the effect of imbalanced annotations on in-context learning.
 In this work, we present the first study on the impact of long-tailed distribution in annotated datasets on in-context learning.
 First, we theoretically show that the class prior distribution affects the prediction of ICL.
-Empirically, we also find that imbalanced class distributions in annotated datasets significantly degrade the performance of ICL across various tasks and selection methods (refer to Figure LABEL:figure:2).
-Moreover, traditional rebalance methods not only fail to ameliorate but, in some cases, exacerbate the issue of class imbalance in ICL.
-This motivates us to develop a universal method that consistently improves ICL’s performance in the presence of imbalanced annotated datasets.
-In this paper, we show that the issue of imbalanced annotated datasets can be mitigated by applying class weighting to the original scoring functions during selection.
-Our method is motivated by accounting for the differences between annotated and test distributions using two-component weights: class-wise weights, which assesses the difference between class prior distributions, and conditional bias, which measures the bias of class-wise weights due to different conditional distributions.
-Therefore, our key idea is to employ the two-component weights to rectify the original scoring functions during selection.
-Specifically, we employ effective number (Cui et al., 2019) as class-wise weights and estimate conditional bias by minimizing the empirical error of a balanced validation set.
-We then employ the two-component weights to modify the original scoring functions during selection.
-In effect, our approach can prevent selecting too many demonstrations from a single class while preserving the effectiveness of the original selection methods.
+Empirically, we also find that imbalanced class distributions in annotated datasets significantly degrade the performance of ICL across various tasks, regardless of selection methods (refer to Figure 1).
+Moreover, traditional rebalancing methods not only fail to ameliorate but, in some cases, exacerbate the issue of class imbalance in ICL.
+This motivates us to develop a universal method that consistently improves ICL’s performance in the presence of imbalanced annotations.
+In this paper, we show that the issue of imbalanced annotations can be mitigated by reweighting the original scores (i.e., cosine similarity score [4]) of imbalanced examples during selection.
+Our method, Reweighting with Conditional Bias (dubbed RCB), is motivated by our analysis of decomposing the distributional differences between imbalanced and test datasets.
+We find that imbalanced datasets affect the performance of ICL through two lenses: class weights – the difference between class prior distributions – and conditional bias, which measures the bias of feature distributions within classes.
+The latter is typically neglected in existing rebalancing methods, leading to suboptimal improvements in ICL performance with imbalanced datasets.
+Therefore, our key idea behind RCB is to incorporate conditional bias into demonstration reweighting during selection.
+Specifically, we calculate the effective number [9] as class weights and estimate conditional bias using Bayesian optimization on a balanced subset.
+Subsequently, a two-component weight containing both class weights and conditional bias is then used to reweight the demonstration scores for selection.
+In effect, our approach prevents over-selection from dominant classes while preserving the efficacy of current selection methods.
 […]
@@
   1. 1.
-    We find that imbalanced class distributions in annotated datasets significantly degrade the ICL’s performance.
-We also show that classical rebalance methods fail to ameliorate the imbalanced problems of ICL.
+    We present a phenomenon: imbalanced annotations significantly degrade the performance of ICL regardless of demonstration selection methods.
+Moreover, classical rebalancing methods fail to address this issue.
   2. 2.
-    We propose a simple and effective method to enhance ICL’s performance with imbalanced annotated datasets.
-The key idea is to estimate the conditional bias and employ the two-component weights to modify the original scoring functions during selection.
+    We propose a simple and complementary method by involving conditional bias to enhance ICL’s performance under class imbalance.
+Our method is computationally efficient and agnostic to demonstration selection methods.
   3. 3.
-    We empirically show that our method can improve the performance of existing selection methods across various imbalance ratios.
-Meanwhile, we also validate the effectiveness of our method on generation tasks.
+    We empirically validate that our methods can improve the ICL performance in both classification and generation tasks across various imbalance ratios.
+Our method can be applied to both open-weight LLMs and APIs, as it only requires access to model outputs.
 2 Preliminary
@@
 input-output pairs) selected from a large annotated dataset with $N$ examples $\mathcal{D}_{c}=\{(\mathrm{\mathbf{x}}_{i},\mathrm{\mathbf{y}}_{i})\}^{N}_{i=1}$.
-Let $f_{\theta}(\mathrm{\mathbf{C}}_{K},\cdot)$ be the ICL model with demonstrations $\mathrm{\mathbf{C}}_{K}$, using the LLM $f$ parameterized by $\boldsymbol{\theta}$.
+Let $f_{\theta}(\mathrm{\mathbf{C}}_{K},\cdot)$ be the ICL model with demonstrations $\mathrm{\mathbf{C}}_{K}$, using the LLM $f$ parameterized by $\bm{\theta}$.
 Given a test input $\mathrm{\mathbf{x}}_{t}$, we generate the output $\mathrm{\mathbf{y}}_{t}$ via the ICL model as:
 $\mathrm{\mathbf{y}}_{t}=f_{\theta}(\mathrm{\mathbf{C}}_{K},\mathrm{\mathbf{x}}_{t})=\mathop{\arg\max}\limits_{\mathrm{\mathbf{y}}}P_{\theta}\left(\mathrm{\mathbf{y}}_{t}|\mathrm{\mathbf{C}}_{K},\mathrm{\mathbf{x}}_{t}\right).$ (1)
-To improve the performance of ICL, previous studies (Liu et al., 2022; Peng et al., 2024; Rubin et al., 2022; Ye et al., 2023a) designed various scoring functions $s(\cdot,\cdot)$ to select demonstrations $\mathrm{\mathbf{C}}_{K}$ from an annotated dataset $\mathcal{D}_{c}$ as:
+To improve the performance of ICL, previous studies [4, 15, 16, 17] designed various scoring functions $s(\cdot,\cdot)$ to select demonstrations $\mathrm{\mathbf{C}}_{K}$ from an annotated dataset $\mathcal{D}_{c}$ as:
 $\mathrm{\mathbf{C}}_{K}=\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right).$ (2)
 […]
-While current selection methods showcase promising performance on common-used benchmarks, their effectiveness may hinge on the distribution of annotation datasets $\mathcal{D}_{c}$.
+While current selection methods showcase promising performance on commonly used benchmarks, their effectiveness may hinge on the distribution of annotation datasets $\mathcal{D}_{c}$.
 For example, ICL might struggle to make accurate predictions for underrepresented groups within these annotated datasets.
@@
 In this work, we begin with the class-imbalance setting of in-context learning in classification tasks [^1] In addition to labels, the imbalance can also occur among various groups, particularly in generation tasks.
-We extend the imbalanced setting to generation tasks in Section 6.
+We extend the imbalanced setting to generation tasks in Section 6.1.
 , where the label space $\mathcal{Y}:=\{1,\ldots,k\}$.
@@
 In particular, the class distribution is such that: $n_{j}\ll n_{k}$, for some $j,k\in\mathcal{Y}$, where $j\neq k$.
-We quantify the imbalance ratio as $\phi=\frac{\min_{j\in\mathcal{Y}}n_{j}}{\max_{j\in\mathcal{Y}}n_{j}}$, and a higher imbalance ratio indicates a more severe class imbalance in the dataset.
+We quantify the imbalance ratio as $\phi=\frac{\max_{j\in\mathcal{Y}}n_{j}}{\min_{j\in\mathcal{Y}}n_{j}}$, and a higher imbalance ratio indicates a more severe class imbalance in the dataset.
 In the real world, class-imbalanced distributions are frequently observed in various datasets.
-For instance, in the Emotion dataset (Saravia et al., 2018), the ‘Joy’ class constitutes 33% of the data, whereas the ‘Surprise’ class makes up only 4%.
-The CodeSearchNet dataset (Husain et al., 2019) includes 29% JavaScript code while Ruby accounts for just 3%, demonstrating the significant imbalance issue among programming languages.
+For instance, in the Emotion dataset [12], the ‘Joy’ class constitutes 33% of the data, whereas the ‘Surprise’ class makes up only 4%.
+The CodeSearchNet dataset [14] includes 29% JavaScript while Ruby accounts for just 3%, demonstrating the significant imbalance issue.
 Therefore, it is crucial to ensure the ICL performance across all classes under the class-imbalanced setting in $\mathcal{D}_{c}$.
@@
 3 Pilot study
-3.1 The effect of imbalanced dataset
-3.1.1 Theoretical Analysis
-We start by presenting Remark 3.1 (Xie et al., 2022), which demonstrates that ICL allows LLMs $f$ parameterized by $\theta$ to learn the context-generated distribution from demonstrations $\mathrm{\mathbf{C}}_{K}$, such that the model’s prediction approximates the output $\mathrm{\mathbf{y}}$ generated by the data generated model $\theta^{*}$ given the test input $\mathrm{\mathbf{x}}$.
-Remark 3.1.
+3.1 The effect of imbalanced data
+Theoretical Analysis
+In this section, we start by presenting Remark 1 [18], which demonstrates that ICL allows LLMs $f$ parameterized by $\theta$ to learn the context-generated distribution from demonstrations $\mathrm{\mathbf{C}}_{K}$, such that the model’s prediction approximates the output $\mathrm{\mathbf{y}}$ generated by the data-generating model $\theta^{*}$ given the test input $\mathrm{\mathbf{x}}$.
+Remark 1
 Assume both demonstrations $\mathrm{\mathbf{C}}_{K}$ and test sample $(\mathrm{\mathbf{x}}_{t},\mathrm{\mathbf{y}}_{t})$ are generated by data generated model $\theta^{*}$.
@@
 $f_{\theta}(\mathrm{\mathbf{C}}_{K},\mathrm{\mathbf{x}}_{t})\approx\mathop{\arg\max}\limits_{\mathrm{\mathbf{y}}}P_{\theta^{*}}(\mathrm{\mathbf{x}}_{t}|\mathrm{\mathbf{y}}_{t})P_{\theta^{*}}(\mathrm{\mathbf{y}}_{t}).$ (3)
-where $P_{\theta^{*}}(\mathrm{\mathbf{x}}_{c})$ and $P_{\theta^{*}}(\mathrm{\mathbf{y}}_{t})$ represent $P_{\theta^{*}}(X=\mathrm{\mathbf{x}}_{t})$ and $P_{\theta^{*}}(Y=\mathrm{\mathbf{y}}_{t})$, respectively.
+where $P_{\theta^{*}}(\mathrm{\mathbf{x}}_{t})$ and $P_{\theta^{*}}(\mathrm{\mathbf{y}}_{t})$ represent $P_{\theta^{*}}(X=\mathrm{\mathbf{x}}_{t})$ and $P_{\theta^{*}}(Y=\mathrm{\mathbf{y}}_{t})$, respectively.
 The Eq.
-3 indicates that the class prior distribution $P(Y)$ affects the prediction of ICL.
-in real-world tasks, the class prior of the test distribution is usually balanced (i.e., a uniform distribution over labels), while the annotated dataset exhibits a long-tailed class distribution.
-Formally, we have $P_{c}(Y=j)\neq P_{t}(Y=j)$ for any class $j\in\mathcal{Y}$.
+(3) indicates that the class prior distribution $P(Y)$ affects the prediction of ICL.
+For the class-imbalanced setting, the class prior of the test distribution is usually balanced, whereas the annotated dataset exhibits a long-tailed class distribution.
 The predicted posterior probability in Eq.
-3 becomes unreliable when there is a large discrepancy of class priors between annotated dataset $P_{c}(Y)$ and test dataset $P_{t}(Y)$.
-3.1.2 Empirical Analysis
-We also conduct experiments on various downstream tasks, including Amazon (Keung et al., 2020), AgNews, Yahoo and Yelp (Zhang et al., 2015), NQ (Kwiatkowski et al., 2019) and CodeSearchNet (Husain et al., 2019)).
-To simulate the class-imbalanced setting, we generate imbalanced datasets with a pre-defined probability (e.g., $\phi$ = 0, 10, 50, 100).
-We evaluate the performance of ICL on a balanced test dataset with three LLMs, including Llama3-8B (AI@Meta, 2024), Mistral-7B (Jiang et al., 2023) and OPT-6.7B (Zhang et al., 2022a).
-Furthermore, we also compare the performance of ICL with various set sizes and selection methods, including Random (Min et al., 2022), TopK (Liu et al., 2022), DPP (Ye et al., 2023a), VoteK (Su et al., 2023), ConE (Peng et al., 2024), ByCS (Wang et al., 2024).
+(3) becomes unreliable when we have $P_{c}(Y=j)\neq P_{t}(Y=j)$ for any class $j\in\mathcal{Y}$.
+Overall, the above theoretical analysis shows that class-imbalanced annotations cause a mismatch between the class prior distributions of the annotated and test datasets, leading to unreliable posterior probability predictions in ICL.
+Our experimental results further validate the analysis.
+Empirical Analysis
+Figure 1: The impact of imbalanced annotations on ICL performance.
+(a) Overall accuracy across four classification tasks.
+(b) Average accuracy for each class in AgNews.
+(c) Overall accuracy across four classification tasks using classical rebalancing methods.
+We also conduct experiments on various downstream tasks, including Amazon [10], AgNews, Yahoo, Yelp [11], NQ [13], and CodeSearchNet [14]).
+To simulate the class-imbalanced setting, we generate imbalanced datasets with a pre-defined probability (e.g., $\phi$ = 1, 10, 50, 100).
+We evaluate the performance of ICL on a balanced test dataset with various LLMs, including OPT-6.7B, -13B, -30B [19], LLAMA-3-8B and -70B [20]; and APIs: ChatGPT-3.5-Turbo [21] and Gemini-2.0-Flash [22].
+We also investigate the performance of ICL with imbalanced annotations using various sizes of demonstrations and selection methods.
 […]
 […]
-Through the above analysis, we find that imbalanced datasets significantly hurt the ICL’s performance.
-Next, we will verify whether employing some classical rebalance methods can mitigate the issue of imbalanced datasets.
-3.2 The failure of classical rebalance methods
-Rebalance methods are widely used to deal with the issue of imbalanced annotated datasets.
-One such approach is to address the bias towards the minor class by adjusting the annotated dataset.
-In this section, we use four different rebalance methods, including over-sampling (Chawla et al., 2002), under-sampling (Liu et al., 2008), stratified sampling (Vilariño et al., 2005) and re-weighting (Cui et al., 2019).
-Specifically, given an annotated dataset with $N$ examples, we select demonstrations as follows:
+3.2 The failure of classical rebalancing methods
+One simple and intuitive approach to deal with the class-imbalanced problem is rebalancing [24].
+In this section, we examine whether four classical rebalancing methods—over-sampling [25], under-sampling [26], stratified sampling [27], and re-weighting [9]—can mitigate the negative effects of imbalanced annotations on ICL.
+Specifically, we select top K examples ranked by the score function $s(\cdot,\cdot)$ from an annotated dataset with $N$ examples as demonstrations:
 $\mathrm{\mathbf{C}}_{K}$ $=\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}^{{}^{\prime}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N^{{}^{\prime}}}\right),$ (4)
-$\mathrm{\mathbf{C}}_{K}$ $=\cup_{i=1}^{k}\operatorname{top}_{\frac{K}{k}}\left(\left\{s(\mathrm{\mathbf{c}}_{j},\mathrm{\mathbf{x}}_{t})\right\}_{j=1}^{n_{i}}\right),$ (5)
+$\mathrm{\mathbf{C}}_{K}$ $=\sum_{i=1}^{k}\operatorname{top}_{\frac{K}{k}}\left(\left\{s(\mathrm{\mathbf{c}}_{j},\mathrm{\mathbf{x}}_{t})\right\}_{j=1}^{n_{i}}\right),$ (5)
 $\mathrm{\mathbf{C}}_{K}$ $=\operatorname{Top}_{K}\left(\left\{w_{i}\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right).$ (6)
 For the over-sampling method in Eq.
-4, we select demonstrations $\mathbf{c}^{{}^{\prime}}_{i}$ from an over-sampling dataset with $N^{{}^{\prime}}$ examples where we repeat the examples of tail classes until their number matches that of head classes.
+(4), we select demonstrations $\mathbf{c}^{{}^{\prime}}_{i}$ from an over-sampling dataset with $N^{{}^{\prime}}$ examples, where we repeat the examples of tail classes until their number matches that of head classes.
 The under-sampling method in Eq.
-4 randomly trims examples belonging to head classes until the number of examples in head classes is equal to that of tail classes.
-Suppose $n_{j}$ represents the size of the class to which $i$-th example belongs, stratified sampling in Eq.
-5 selects $\frac{K}{k}$ demonstrations from each class with $n_{j}$ examples.
+(4) randomly trims examples belonging to head classes until the number of examples in head classes is equal to that of tail classes.
+Suppose $n_{j}$ represents the size of the class to which the $i$-th example belongs, stratified sampling in Eq.
+(5) selects $\frac{K}{k}$ demonstrations from each class with $n_{j}$ examples.
 For the re-weighting method in Eq.
-6, according to the classical reweighing method for fine-tuning (Cui et al., 2019), we select the top $K$ examples based on the scoring functions $s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})$ multiplied by the class-wise weights $w_{j}$ where $w_{j}=(1-\alpha)/(1-\alpha^{n_{j}})$ and $\alpha=(N-1)/N$.
-We conduct experiments to verify the effectiveness of these rebalance strategies across various downstream tasks within the framework of ICL.
-The results in Figure 3 show that employing over-sampling and under-sampling can improve ICL’s performance when annotated datasets exhibit long-tail distributions.
-However, these rebalance strategies can only achieve limited effects for most downstream tasks, indicating poor generalization to ICL.
-Specifically, the limited effect of the oversampling method (Chawla et al., 2002) suggests that repeated examples do not provide additional information to LLMs to enhance the performance of ICL.
-Under-sampling (Liu et al., 2008) may remove key information of head classes.
-Notably, the failure of the stratified sampling method in ICL is mainly due to the decreased performance of selection methods such as TopK (Liu et al., 2022) and DPP (Ye et al., 2023a).
-The detailed results of generation tasks have been reported in Section 6.1.
-Our empirical analysis shows that existing rebalance methods cannot effectively address the issue of imbalanced datasets for ICL.
-This motivates us to design new rebalance methods for ICL that can universally enhance the performance of existing selection methods.
-Figure 3: Comparison of performances on four classification tasks with different imbalance ratios using four different rebalance methods.
-We report the average accuracy (%) of six existing selection methods ($K=8$) using OPT-6.7B (Zhang et al., 2022a).
+(6), we select the top $K$ examples based on the scoring functions $s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})$ multiplied by the class weights $w_{j}$ where $w_{j}=(1-\alpha)/(1-\alpha^{n_{j}})$ and $\alpha=(N-1)/N$ [9].
+Figure 1 (c) shows that the rebalancing methods can only achieve limited effects for most downstream tasks, indicating poor generalization to ICL.
+Intuitively, oversampling uses repeated examples that may not provide additional information to LLMs for ICL performance, while undersampling may remove key information of head classes.
+For instance, with an imbalanced ratio of 100, over-sampling boosts the performance of ICL from 45.10 to 46.30, yielding limited improvement.
+In contrast, stratified sampling, which selects demonstrations equally from each class, reduces the performance of ICL from 45.10 to 42.46, as it undermines the effectiveness of high-performing selection methods like TopK [4].
+Notably, we provide an in-depth understanding of the failure through the decomposition of the distributional differences in Subsection 4.1.
 4 Methodology
-In this section, we first decompose the distributional differences between annotated and test datasets to explain why classical rebalance methods fail for ICL.
-In light of this, we propose a novel method to enhance the performance of ICL with imbalanced annotated datasets.
+Figure 2: The Kullback-Leibler (KL) divergence between $P_{c}(X|Y)$ and $P_{t}(X|Y)$ for each class in the AgNews (a), Amazon (b) and Yelp (c) datasets.
+The KL divergence validates the existence of conditional bias in imbalanced datasets.
 4.1 Decomposition of the distributional differences
-We aim to design a suitable demonstration selection method, which is used to select demonstrations for the given test input $\mathrm{\mathbf{x}}_{t}$, to minimize expected $\operatorname{Error}$ value across all classes:
-$\operatorname{Error}=\mathbb{E}_{P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right],$
+For class-imbalanced annotated datasets $\mathcal{D}_{c}$, our goal is to select suitable demonstrations for a given test input $\mathrm{\mathbf{x}}_{t}$ in order to minimize the expected error $\operatorname{Error}$ across all classes:
+$\operatorname{Error}=\mathbb{E}_{P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right],$ (7)
 where $M[\cdot,\cdot]$ measures the mismatched degree between real output and the generated output conditioned on demonstrations $\mathrm{\mathbf{C}}_{K}$ and test input $\mathrm{\mathbf{x}}_{t}$.
-For example, for classification tasks, $M[\cdot,\cdot]$ denotes error rate.
-For question answering tasks, $M[\cdot,\cdot]$ represents the negative value of EM score.
-We apply the importance sampling trick to connect the expected $\operatorname{Error}$ and decompose the distributional differences into two-component weights as follows:
-$\mathbb{E}_{P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right],$
-$=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right]\frac{P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})},$
-$=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right](\bm{w}+\bm{\beta}).$
-where $P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})/P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})$ measures the distributional difference, $\bm{w}=P_{t}(Y)/P_{c}(Y)$ denotes class-wise weights and $\bm{\beta}=\frac{P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{y}})}\left(\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1\right)$ is conditional bias.
-The detailed proof can be found in Appendix A.2.
-Notably, these existing rebalance methods (i.e.
-re-weighting (Cui et al., 2019)) assume a shared conditional distribution between annotated and test datasets $P_{c}(X|Y)=P_{t}(X|Y)$ and only focus on how to use the class-wise weights $\bm{w}$ to process imbalanced datasets.
-However, in the real world, the conditional distribution $P(X|Y)$ is also prone to changes from the annotated dataset to the test dataset due to target shift and concept shift, especially in imbalanced datasets (Gopalan et al., 2013; Jamal et al., 2020; Gu et al., 2022).
-We also conduct experiments to verify whether there is a shared conditional distribution between annotated and test datasets $P_{c}(X|Y)=P_{t}(X|Y)$ in imbalanced datasets.
-Specifically, for each class, we firstly transform text input from each class into bag-of-words representation, employ the Latent Dirichlet Allocation (LDA) model to derive the document-topic distributions and then compute Kullback-Leibler (KL) divergence between annotated conditional distribution $P_{c}(X|Y)$ and test conditional distribution $P_{t}(X|Y)$.
-For example, as shown in Figure LABEL:figure:4, the KL divergence between annotated and test datasets is unequal to zero and the assumption that $\beta=0$ is difficult to be satisfied.
-In this paper, we relax the above assumption and allow a different conditional distribution of annotated and test datasets (e.g., $P_{c}(X|Y)\neq P_{t}(X|Y)$).
+For example, $M[\cdot,\cdot]$ is the error rate for classification tasks and the negative value of the EM score for question-answering tasks.
+Next, we apply the importance sampling trick to connect the expected $\operatorname{Error}$ of ICL with the annotated dataset to test dataset as follows:
+$\operatorname{Error}$ $=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right]\frac{P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})},$
+$=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right](\bm{w}+\bm{\beta}).$ (8)
+where $P_{t}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})/P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})$ measures the distributional difference, $\bm{w}=P_{t}(Y)/P_{c}(Y)$ denotes class weights and $\bm{\beta}=\frac{P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{y}})}\left(\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1\right)$ is conditional bias.
+The proof is shown in Appendix B.
+For an imbalanced dataset, it is easily observed that there is a large discrepancy between the class priors of the annotated and test distributions, resulting in $\bm{w}\neq 1$.
+Due to target shift and concept shift caused by limited sampling in the tailed classes, the conditional distribution between the annotated and test datasets usually shows significant differences, thereby $\bm{\beta}\neq 0$.
+To verify the existence of conditional bias, we employ the Latent Dirichlet Allocation (LDA) model to derive the document-topic distributions for each class and then compute the Kullback-Leibler (KL) of conditional distribution between the annotated and test datasets.
+Figure 2 shows that the KL divergence of conditional distribution between the annotated and test datasets is not equal to zero, thereby verifying the existence of conditional bias $\bm{\beta}$ in the imbalanced datasets.
+Overall, the above analysis illustrates that the performance of ICL is affected by both class weights $\bm{w}$ and conditional bias $\bm{\beta}$; however, the latter is typically neglected in existing rebalancing methods in Subsection 3.2, leading to suboptimal performance of ICL with imbalanced annotations.
+4.2 Reweighting with Conditional Bias
+The previous analysis shows that the performance of ICL significantly deteriorates when demonstrations are selected from an imbalanced dataset without considering the conditional bias $\bm{\beta}$.
 However, directly computing the conditional bias $\bm{\beta}$ is non-trivial as real conditional distributions of annotated and test datasets are unknown.
-To circumvent the issue, we aim to design an effective method to estimate the conditional bias $\bm{\beta}$.
-4.2 Modeling the conditional bias $\bm{\beta}$
-Motivated by the previous analysis, we employ two-component weights, including class-wise weights $\bm{w}$ and conditional bias $\bm{\beta}$, to modify the original scoring functions during selection.
-We let the class-wise weights resemble the empirically successful design in the literature.
-Our key idea is to estimate conditional bias $\bm{\beta}$ by achieving empirical $\operatorname{Error}$ minimization of a balanced validation set.
-Constructing balanced validation set
-Given an imbalanced dataset $\mathcal{D}_{c}$ with $N$ examples, the balanced validation set $\mathcal{D}_{b}$ is constructed by randomly selecting $n_{b}$ examples from $j$-class dataset $\mathcal{D}_{j}$ as follows:
-$\mathcal{D}_{b}=\cup_{j=1}^{k}\left\{\operatorname{RandomSample}(\mathcal{D}_{j},n_{b})\right\},$ (7)
-where $\operatorname{RandomSample}(\mathcal{D}_{j},n_{b})$ denotes randomly selecting $n_{b}$ examples from each class set $\mathcal{D}_{j}$.
-Notably, $n_{b}$ is constrained by the condition $n_{b}<\min(n_{1},...,n_{k})$.
-This process results in a validation set that is balanced across all classes, with each class contributing the same number of examples, similar to the composition of the test dataset.
-Meanwhile, $\mathcal{D}_{r}$ denotes the imbalanced remaining dataset.
+In what follows, we propose Reweighting with Conditional Bias (RCB), a complementary strategy to enhance the ICL performance under class imbalance.
+Our key idea is to incorporate conditional bias into the demonstration reweighting in selection.
+With this in mind, we present the details of our approach in the following.
+Selecting a balanced subset
+Given an imbalanced dataset $\mathcal{D}_{c}$ with $n_{j}$ examples for the $j$-th class, we select a balanced subset $\mathcal{D}_{b}$ by uniformly sampling $n_{b}$ examples from per class of $\mathcal{D}_{c}$:
+$\mathcal{D}_{b}=\operatorname{UniformSample}(\mathcal{D}_{c},n_{b}),$ (9)
+where $\operatorname{UniformSample}(\mathcal{D}_{c},n_{b})$ denotes uniformly selecting $n_{b}$ examples per class of $\mathcal{D}_{c}$, $n_{b}$ is constrained by the condition $n_{b}<\min_{j\in\mathcal{Y}}n_{j}$ to ensure $n_{b}$ does not exceed the size of the smallest class.
+Through the selection, we construct a balanced subset where all classes contribute an equal number of examples, aligning with the composition of the test dataset.
+It is worth noting that the selected balanced subset can be also used.
+Estimating the conditional bias
+We let the class weights $\bm{w}$ resemble the empirically successful design from previous literature [9, 28].
+We continue to use effective numbers [9] as class weights $\bm{w}$, defined as $w_{j}=(1-\alpha)/(1-\alpha^{n_{j}})$ where $\alpha=(N-1)/N$ and $N$ denotes the total number of examples in imbalanced dataset $\mathcal{D}_{c}$.
+In practice, since the true conditional distribution cannot be directly computed, we approximate the conditional bias $\bm{\beta}$ by leveraging the balanced subset $\mathcal{D}_{b}$:
+$\frac{1}{|\mathcal{D}_{b}|}\sum_{i=1}^{|\mathcal{D}_{b}|}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\{(\bm{w}+\bm{\beta})\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{b})\right\}_{i=1}^{|\mathcal{D}_{r}|}),\mathrm{\mathbf{x}}_{b}\right),\mathrm{\mathbf{y}}_{b}\right],$ (10)
+where we compute the mismatched degree $M$ of the balanced subset $\mathcal{D}_{b}$ between real output $\bm{y}_{b}$ and ICL’s output conditioned on demonstrations selected from a remaining imbalanced dataset $\mathcal{D}_{r}=\mathcal{D}_{c}\setminus\mathcal{D}_{b}$, while preserving the original imbalance.
+We employ $\operatorname{Top}_{K}$ to select $K$ highest-ranked demonstrations based on reweighting existing demonstration scores for selection $s(\cdot,\cdot)$ via $\bm{w}+\bm{\beta}$.
+We use a Bayesian optimization framework, which is suitable for non-differentiable and black-box function optimization, to estimate conditional bias $\bm{\beta}$ with the lowest mismatched degree on the balanced subset.
+We show the detailed bayesian optimization process in Appendix C.
+In this way, we can estimate conditional bias $\bm{\beta}$ to measure the bias of feature distributions within classes.
+Re-weighting with class weights $\bm{w}$ and conditional bias $\bm{\beta}$
+Given a test input $\mathrm{\mathbf{x}}_{t}$, we first select $K^{\prime}$ candidates from the imbalanced dataset $\mathcal{D}_{c}$ using existing selection methods.
+A large $K^{\prime}$ ensures that all classes (especially tailed classes) are included in the subset of candidates selected.
+For each candidate $\mathbf{c}_{i}$, we use $\bm{w}+\bm{\beta}$ to reweight its original score $s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})$ and employ $\operatorname{argsort}$ to sort the adjusted score in ascending order, producing sorted indices $\mathcal{I}$:
+$\mathcal{I}=\operatorname{argsort}\left\{\right(\bm{w}+\bm{\beta}\left)\times s\right(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t}\left)\right\}_{i=1}^{K^{{}^{\prime}}},$
+We select these candidates with $K$ low-ranking indices in the sorted list $\mathcal{I}$ as the final demonstrations:
+$g(\mathrm{\mathbf{c}}_{i})=\mathbbm{1}\left(\operatorname{Loc}(\mathrm{\mathbf{c}}_{i},\mathcal{I})\leq K\right).$
+where $\mathbbm{1}$ is the indicator function and $\operatorname{Loc}(\mathrm{\mathbf{c}}_{i},\mathcal{I})$ return the index of $\mathrm{\mathbf{c}}_{i}$ in the sorted list $\mathcal{I}$.
+After the above steps, we establish the final $K$ demonstration set for in-context learning.
+Noticeably, our method offers several compelling advantages:
+  • Algorithm-agnostic: Our method can be easily incorporated into existing demonstration selection methods, improving the performance of ICL with imbalanced annotations.
+  • Easy to use: Our method requires access only to the model outputs and integrates effortlessly with any LLMs (see Figure 3 (d)).
+Our method is insensitive to the size of the balance subset $|\mathcal{D}_{b}|$ and class weights $\bm{w}$ (see Figure 3 (a) and (b)).
+  • High efficiency: Our method only estimates the conditional bias for each class once for an imbalanced dataset.
+Thus, our method introduces negligible computational costs compared to in-context learning.
+5 Experiments
+5.1 Experimental Setup
+Datasets and Evaluation For our evaluations, we verify the effectiveness of our method on four benchmark datasets, including Sentiment Classification (Amazon [10], Yelp [11]) and Topic Classification (AgNews, Yahoo [11]).
+Due to limited space, these tasks’ input/output, statistics and split are reported in the Appendix D.
+To simulate the issue of imbalanced data, we generate imbalanced datasets with pre-defined imbalance ratios $\phi$ (e.g., 1, 10, 50, 100) and we evaluate the performance of ICL on another corresponding class-balanced test dataset [24, 28].
+We report average accuracy (%) with standard deviation (over 3 runs).
+Models and Baselines We utilize various large language models, including open-weight models: OPT-6.7B, OPT-13B, OPT-30B, LLAMA-3-8B and LLAMA-3-70B; and APIs: ChatGPT-3.5-Turbo and Gemini-2.0-Flash.
+We use Bert-base-uncased sentence encoder [29] as the similarity tokenizer.
+We also conduct experiments with existing selection methods as baselines, including Random [30], TopK [4], DPP [17], VoteK [31], ConE [15], ByCS [6].
+For hyperparameters, we set the size of the balanced subset as $|\mathcal{D}_{b}|=100$ and $K^{\prime}=1600$ by default.
+The details of our experiments are presented in Appendix D.
 […]
 $K$ 8 16
-Imbalance Ratio 0 10 50 100 0 10 50 100
-Random 32.63$\pm$0.73 32.26$\pm$1.32 30.91$\pm$1.20 30.21$\pm$0.83 35.25$\pm$0.46 36.17$\pm$0.91 33.24$\pm$1.24 32.51$\pm$0.65
-+Ours 32.44$\pm$0.92 36.66$\pm$1.05 36.45$\pm$1.26 37.63$\pm$1.79 34.68$\pm$0.53 36.12$\pm$0.99 38.74$\pm$1.36 40.32$\pm$1.79
-TopK 49.90$\pm$0.05 46.51$\pm$0.76 41.51$\pm$1.11 39.62$\pm$1.22 52.31$\pm$0.09 48.18$\pm$0.98 43.34$\pm$0.91 40.95$\pm$0.61
-+Ours 49.75$\pm$0.41 49.62$\pm$0.62 46.48$\pm$1.07 44.44$\pm$1.40 51.73$\pm$0.49 51.19$\pm$1.08 48.21$\pm$1.01 45.52$\pm$1.40
-DPP 50.56$\pm$0.06 46.59$\pm$0.79 42.54$\pm$1.09 40.26$\pm$1.22 53.66$\pm$0.15 48.38$\pm$0.98 43.53$\pm$0.91 41.83$\pm$0.64
-+Ours 50.36$\pm$0.31 49.47$\pm$1.57 46.50$\pm$1.26 45.16$\pm$1.64 52.84$\pm$0.27 50.97$\pm$1.14 48.16$\pm$1.18 46.77$\pm$1.48
-VoteK 48.69$\pm$0.14 45.71$\pm$0.88 41.51$\pm$1.13 39.94$\pm$0.95 52.62$\pm$0.24 48.02$\pm$0.71 42.72$\pm$0.91 40.57$\pm$0.60
-+Ours 48.57$\pm$0.39 48.21$\pm$0.97 46.39$\pm$1.23 44.77$\pm$1.27 52.62$\pm$0.44 50.97$\pm$0.98 48.05$\pm$0.82 45.84$\pm$1.05
-ConE 48.08$\pm$0.86 44.55$\pm$1.51 40.35$\pm$1.29 38.03$\pm$1.31 51.18$\pm$0.58 46.09$\pm$0.79 41.30$\pm$0.63 38.92$\pm$0.72
-+Ours 48.51$\pm$0.70 46.89$\pm$0.85 45.09$\pm$1.08 43.21$\pm$1.32 50.95$\pm$0.46 48.32$\pm$0.84 46.83$\pm$1.07 44.13$\pm$0.79
-ByDC 50.07$\pm$0.45 45.00$\pm$0.78 41.10$\pm$1.23 38.90$\pm$1.36 53.15$\pm$0.44 46.16$\pm$0.75 41.90$\pm$0.76 38.92$\pm$0.62
-+Ours 49.82$\pm$0.39 48.54$\pm$0.94 46.60$\pm$1.11 44.52$\pm$1.05 52.61$\pm$0.55 49.98$\pm$1.00 47.33$\pm$1.23 44.39$\pm$1.14
-Bayesian optimizing conditional bias
-In practice, since the data distribution is unknown, we use empirical $\operatorname{Error}$ minimization of the balanced validation set $\mathcal{D}_{b}$ to achieve an empirical estimate of the underlying data distribution.
-Typically, we minimize the empirical $\operatorname{Error}$ as follows:
-$\frac{1}{|\mathcal{D}_{b}|}\sum_{i=1}^{|\mathcal{D}_{b}|}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\{(w_{i}+\beta)\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{b})\right\}_{i=1}^{|\mathcal{D}_{r}|}),\mathrm{\mathbf{x}}_{b}\right),\mathrm{\mathbf{y}}_{b}\right].$
-where we still use effective numbers (Cui et al., 2019) as class-wise weights where $w_{i}=(1-\alpha)/(1-\alpha^{n_{j}})$ and $\alpha=(N-1)/N$.
-We estimate the conditional bias $\beta$ by employing a Bayesian optimizing framework, which is suitable for non-differentiable and black-box function optimization.
-In this way, the estimated conditional bias $\beta$ can be regarded as an approximation of the conditional bias between annotated and test datasets.
-We present our overall optimizing process in Appendix A.3.
-Re-weighting with adjusted class-wise weights
-We use existing selection methods to select a subset of candidates (e.g.
-$K^{{}^{\prime}}$=100).
-We sort all candidates in increasing order by $(w_{i}+\beta)\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})$ and obtain the indices for the sorted scores as $\mathcal{I}$.
-And we select these candidates with low-ranking examples in the sorted list $\mathcal{I}$:
-$g(\mathrm{\mathbf{c}}_{i})=\mathrm{1}\left(\operatorname{Loc}(\mathrm{\mathbf{c}}_{i},\mathcal{I})\leq K\right),$ (8)
-$\mathcal{I}=\operatorname{argsort}\left\{(w_{i}+\beta)s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t}))\right\}_{i=1}^{K^{{}^{\prime}}}.$ (9)
+Imbalance Ratio 1 10 50 100 1 10 50 100
+Random 47.70$\pm$0.77 43.72$\pm$1.26 39.59$\pm$0.92 38.01$\pm$0.58 50.68$\pm$0.49 46.65$\pm$0.72 41.30$\pm$0.99 39.90$\pm$0.54
++Ours 47.62$\pm$1.06 47.81$\pm$1.40 46.51$\pm$0.90 47.50$\pm$1.13 50.11$\pm$0.70 50.50$\pm$0.60 50.41$\pm$1.35 51.01$\pm$2.19
+TopK 58.32$\pm$0.03 54.01$\pm$0.68 48.52$\pm$0.99 46.78$\pm$1.50 59.93$\pm$0.05 55.57$\pm$0.58 50.78$\pm$0.88 48.65$\pm$0.88
++Ours 58.12$\pm$0.65 57.46$\pm$0.50 54.14$\pm$1.31 51.92$\pm$1.22 59.53$\pm$0.37 58.94$\pm$0.93 55.58$\pm$0.75 52.83$\pm$1.32
+DPP 58.50$\pm$0.00 53.90$\pm$0.79 49.37$\pm$0.78 47.25$\pm$1.11 60.00$\pm$0.07 55.54$\pm$0.85 51.30$\pm$0.74 49.58$\pm$0.58
++Ours 58.20$\pm$0.34 56.76$\pm$1.09 53.25$\pm$1.32 51.93$\pm$1.13 59.90$\pm$0.11 58.40$\pm$0.94 55.06$\pm$1.10 53.34$\pm$1.26
+VoteK 58.07$\pm$0.14 53.51$\pm$0.78 48.75$\pm$0.92 47.21$\pm$1.26 60.00$\pm$0.15 55.53$\pm$0.49 50.46$\pm$0.79 48.42$\pm$0.90
++Ours 57.92$\pm$0.30 56.20$\pm$0.68 53.95$\pm$1.34 51.84$\pm$1.01 59.89$\pm$0.22 58.23$\pm$0.81 55.45$\pm$0.63 52.93$\pm$0.89
+ConE 57.36$\pm$0.51 52.41$\pm$1.09 47.86$\pm$1.33 45.58$\pm$1.33 58.60$\pm$0.51 53.97$\pm$0.71 49.25$\pm$0.70 46.87$\pm$0.80
++Ours 57.59$\pm$0.62 54.75$\pm$0.91 52.45$\pm$0.70 50.12$\pm$1.29 58.85$\pm$0.41 56.06$\pm$0.81 54.19$\pm$0.74 51.17$\pm$0.75
+ByDC 58.38$\pm$0.40 52.62$\pm$0.87 47.95$\pm$1.09 45.78$\pm$1.17 60.04$\pm$0.31 53.91$\pm$0.54 49.66$\pm$0.77 47.01$\pm$0.71
++Ours 58.16$\pm$0.37 57.24$\pm$0.78 53.25$\pm$1.27 51.45$\pm$0.93 60.03$\pm$0.28 57.53$\pm$0.89 54.53$\pm$0.70 51.44$\pm$0.79
+Average 56.39$\pm$0.31 51.69$\pm$0.91 47.01$\pm$1.00 45.10$\pm$1.16 58.21$\pm$0.26 53.53$\pm$0.65 48.79$\pm$0.82 46.74$\pm$0.73
+56.27$\pm$0.56 55.04$\pm$0.89 52.26$\pm$1.14 50.79$\pm$1.12 58.05$\pm$0.35 56.61$\pm$0.83 54.08$\pm$0.88 52.16$\pm$1.20
+5.2 Main Results
 […]
-  • Flexible: Our method can be easily incorporated into existing selection methods under the class-imbalanced setting in $\mathcal{D}_{c}$.
-  • Easy to use: Our method does not require heavy hyperparameter tuning, as it is insensitive to balance set $\mathcal{D}_{b}$ or class-wise weights $\bm{w}$ (see Figure LABEL:figure:5 (a) and (b)).
-  • Model-agnostic: Our method requires access only to the model outputs and integrates effortlessly with any LLMs (see Figure LABEL:figure:5 (c) and (d)).
-5 Experiments
-5.1 Experimental Setup
-Datasets and Evaluation For our evaluations, we verify the effectiveness of our method on four benchmark datasets, including Sentiment Classification (Amazon (Keung et al., 2020), Yelp (Zhang et al., 2015)) and Topic Classification (AgNews, Yahoo (Zhang et al., 2015)).
-Due to limited space, these tasks’ input/output, statistics and split are reported in the Appendix A.4.
-To simulate the issue of imbalance data, we generate imbalanced datasetss with pre-defined imbalance ratios $\phi$ (e.g., 0, 10, 50, 100) and we evaluate the performance of ICL on another corresponding class-balanced test dataset.
-We report average accuracy (%) with standard deviation (over 3 runs).
-Models and Baselines We utilize various large language models, including Llama3-8B (AI@Meta, 2024), Mistral-7B (Jiang et al., 2023) and OPT-6.7B (Zhang et al., 2022a).
-We use Bert-base-uncased sentence encoder (Devlin et al., 2019) as the similarity tokenizer.
-We also conduct experiments with existing selection methods as baselines, including Random (Min et al., 2022), TopK (Liu et al., 2022), DPP (Ye et al., 2023a), VoteK (Su et al., 2023), ConE (Peng et al., 2024), ByCS (Wang et al., 2024).
-For hyperparameters, we set the size of the balanced validation set as $|\mathcal{D}_{b}|=100$ by default.
-The details of our experiments are presented in Appendix A.4.
-5.2 Main Results
-Can our method improve the performance of existing selection methods?
-Table 1 presents the average test accuracy of in-context learning with different selection methods on four benchmark datasets using OPT-6.7B, under various imbalance ratios.
-A salient observation is that our method drastically improves the performance of ICL with imbalanced annotated datasets.
-For example, for the 100 imbalance ratio, our method improves the average accuracy of six existing selection methods from 37.93 to 43.29 – a 5.46 of direct improvement.
-More importantly, we show that our method can boost performance for a wide range of selection methods, such as TopK and DPP.
-For example, we observe that, for the 100 imbalance ratio, the test accuracy of the TopK is improved to 4.57 when integrating our methods.
-Table 1 also presents the results for different demonstration numbers $K$, suggesting that our method can consistently improve the performance of ICL against imbalanced datasets with varying demonstration numbers $K$.
-The detailed results are reported in Appendix A.6.3.
-How does the balanced validation set affect the robustness of our method?
-In Figure LABEL:figure:5 (a), we ablate how the size of balanced validation set $\mathcal{D}_{b}$ affects the effectiveness of our method (cf.
+How does the balanced subset affect the robustness of our method?
+In Figure 3 (a), we examine how the size of the balanced subset $\mathcal{D}_{b}$ affects the effectiveness of our method (cf.
 Eq.
-7).
-The baselines indicates all candidate demonstrations are selected without our method.
-It’s noteworthy that our method shows robustness to the choice of the size of the balanced validation set $|\mathcal{D}_{b}|$.
-Even when we set $|\mathcal{D}_{b}|=50$, it still yields significant improvements in ICL performance on four classification tasks across six selection methods against imbalanced datasets.
-Due to space constraints, we only report the average results of multiple baselines on various downstream tasks.
-Is our method effective with different class-wise weights $\bm{w}$?
-In the paper, we employ effective numbers as class-wise weights $\bm{w}$, which have been confirmed to empirically successfully deal with imbalance problems in previous studies (Cui et al., 2019; Jamal et al., 2020).
-We also verify whether the type of class-wise weights affects the performance of our methods.
-Specifically, we employ class frequency $w_{j}=\frac{n_{j}}{N}$ (Kang et al., 2020; Shi et al., 2023) as an alternative class-wise weights to verify the effectiveness of our method.
-Our results in Figure LABEL:figure:5 (b) show that our method can improve the ICL’s performance across various imbalanced ratios and achieve similar accuracy to the effective numbers.
-Is our method effective with different model architectures?
-To demonstrate that our proposed method is model-agnostic, we conduct experiments on diverse open-sourced LLMs (Llama3-8B (AI@Meta, 2024) and Mistral-7B (Jiang et al., 2023)) and present the average results of four common benchmarks using the two LLMs in Figure LABEL:figure:5 (c).
-From the results, we observe that our method consistently improves test performance on various downstream tasks when using different model architectures.
-For instance, using our method boosts the test accuracy from 47.10 to 54.31, a 7.21 of direct improvement with a 100 imbalance ratio.
-Therefore, our method can improve the performance of ICL using various open-source LLMs against imbalanced annotated datasets.
-More details are presented in Appendix A.6.4.
-Is our method effective with closed-sourced LLMs?
-While our method has demonstrated strong promise in open-source LLMs, we also verify its effectiveness using closed-source models.
-Specifically, we tested our method on ChatGPT-3.5-Turbo across four selection methods (Random, TopK, DPP and VoteK) with various imbalance ratios.
-Figure LABEL:figure:5 (d) reports the average results for ChatGPT-3.5-Turbo, showing that imbalanced annotated datasets significantly hurt the ICL performance of closed-source LLMs when employing these existing selection methods.
-After integrating our method, Figure LABEL:figure:5 (d) also shows significant improvements in average test accuracy across four downstream tasks, indicating the effectiveness of our method in addressing imbalances in closed-source LLMs.
-Can our method improve real-world imbalanced dataset We verify the effectiveness of our method on a real-world imbalanced dataset.
-The Emotion (Saravia et al., 2018) dataset has a long-tailed distribution (refer to Figure 1).
-Figure 18 shown that our method consistently improves existing selection methods and outperforms existing rebalance methods, including over-sampling (Chawla et al., 2002), under-sampling (Liu et al., 2008), stratified sampling (Vilariño et al., 2005), and re-weighting (Cui et al., 2019).
-For example, with OPT-6.7B (Zhang et al., 2022a), using our method boosts the average accuracy of six selection methods from 21.22 to 23.36, a relative improvement of 10.08%.
+9).
+The Vanilla indicates all candidate demonstrations are selected without our method.
+It’s noteworthy that our method shows robustness to the choice of the size of the balanced subset $|\mathcal{D}_{b}|$.
+Even when we set $|\mathcal{D}_{b}|=50$, it still yields significant improvements in ICL performance on AgNews and Yahoo datasets across six selection methods against imbalanced datasets.
+Notably, the balanced subset can be reused in the process of demonstration selection.
+Thus, our method is applicable in scenarios with extremely limited data from tail classes.
+Is our method robust with different class weights $\bm{w}$?
+In the paper, we employ effective numbers as class weights $\bm{w}$, which have been confirmed to successfully deal with imbalance problems in previous studies [9, 28].
+We employ class frequency $w_{j}=\frac{n_{j}}{N}$ [24, 32] as an alternative class weights to verify whether the type of class weights affects the performance of our methods.
+Figure 3 (b) shows that our method with class frequency can improve the ICL’s performance across various imbalance ratios and achieve accuracy similar to that of effective numbers.
+Does our method work with the different number selected candidates $K^{\prime}$?
+We set a large value of $K^{\prime}$ to ensure that all classes (especially tailed classes) are included in the subset of candidates selected by existing methods.
+Given a small $K^{\prime}$, the candidate subset may contain too few examples from tailed classes (even missing), leading to suboptimal performance.
+Therefore, the $K^{\prime}$ should be larger in cases of larger imbalanced ratios.
+We validate the claim by conducting experiments with varying $K^{\prime}$ on datasets with various imbalanced ratios.
+The result in Figure 3 (c) demonstrates that one should set a sufficiently large $K^{\prime}$ to achieve the best performance.
+Is our method effective with different model architectures and sizes?
+To show our proposed method is model-agnostic, we conduct experiments on a diverse collection of model architectures and sizes, including open-weight models: OPT-6.7B, OPT-13B, OPT-30B, LLAMA-3-8B and LLAMA-3-70B; and APIs: ChatGPT-3.5-Turbo and Gemini-2.0-Flash.
+The results (vanilla/+ours) in the Figure 3 (d) present the same phenomenon as the main experiments in the manuscript: the ICL performance of LLMs gets worse at larger imbalanced ratios, and our method can significantly improve the performance, especially at large imbalanced ratios (e.g., 100).
+For instance, for LLAMA-3-8B [20], using our method boosts the average accuracy from 65.72 to 73.20, a 7.48 of direct improvement with a 100 imbalance ratio.
+The detailed results of each model are presented in the Table 2.
+Can our method work in real-world imbalanced datasets?
+We verify the effectiveness of our method on a real-world imbalanced dataset.
+In Emotion [12], a few sentiments make a large contribution and data tend to show a long-tailed distribution (refer to Figure 5)..
+For example, “joy“ and “sadness“ are head classes, while most other classes, such as “love“ and “surprise“, are tail classes.
+Table 3 shows that our method consistently improves existing selection methods and outperforms existing rebalancing methods.
+For example, with OPT-6.7B [19], using our method boosts the average Macro-F1 of six selection methods from 32.12 to 36.69, a direct improvement of 4.57.
 These results verify that our method is effective in improving ICL’s performance in real-world imbalanced scenarios.
+Figure 3: Results for ablation studies.
+(a) Ablation on the different balanced subset sizes.
+(b) Ablation on the different class weights (c) Ablation on the different selected candidates.
+(d) Ablation on the different model architectures and sizes.
 6 Discussion
@@
 […]
-Table 2 demonstrates that imbalanced annotated dataset significantly hurt ICL’s performance on generation tasks.
-Additionally, Table 2 also confirms the ineffectiveness of classical data balancing methods, including over-sampling (Chawla et al., 2002), under-sampling (Liu et al., 2008), stratified sampling (Vilariño et al., 2005), and re-weighting (Cui et al., 2019).
-After integrating our method, the existing selection methods show significant improvements in inference performance, indicating the generalization of our method to text generation.
-For example, our method improves the average exact match score from 21.20 to 22.93, a notable improvement of 0.73 over six existing demonstration methods.
-The detailed results of CodeSearchNet (Husain et al., 2019) task are presented in Table 7.
-Table 2: The average EM score of NQ dataset across six selection methods with various imbalance ratios. Bold numbers are the best results.
-Imbalance Ratio 0 10 50 100
-Baselines 23.67 22.47 21.33 21.20
-+Over-Sampling 23.60 22.53 22.53 22.33
-+Under-Sampling 23.60 19.87 18.47 17.80
-+Stratified-Sampling 22.15 20.60 18.82 18.47
-+Re-Weighting 23.33 22.60 20.47 20.20
-+Ours 24.00 23.87 23.00 22.93
+[…]
+Figure 4: The impact of imbalanced annotations on the performance of In-Context Learning (ICL) in generation tasks, including Natural Questions (NQ) (a) and CodeSearchNet (b).
+Figure (c) shows the result of imbalanced datasets with an increasing number of examples on four classification tasks.
 6.2 Does a larger dataset address the imbalance issue?
-Table 3: Average classification accuracy (%) of Baselines / +Ours methods on the Yahoo dataset with various imbalance ratios. Bold numbers are superior results.
-Imbalance Ratio 10 50 100
-Baseline/+Ours
-Random 37.87/38.07 36.20/39.87 33.13/37.73
-TopK 47.33/48.87 45.00/47.67 43.07/47.87
-DPP 49.60/50.27 48.00/49.33 44.47/47.80
-VoteK 47.13/47.80 43.87/47.47 44.67/47.73
-ConE 47.53/49.13 45.27/46.80 41.47/45.53
-ByDC 48.40/49.33 44.27/47.47 42.17/46.50
-Average 46.64/47.25 43.77/46.44 41.50/45.53
-While our analysis has shown that imbalanced annotated datasets negatively impact in-context learning by reducing the number of examples in tail classes, one might wonder if this negative effect is due to the reduced size of the dataset rather than a shift in class prior distribution.
-To address this, we verify the negative impact of imbalanced annotated datasets and the effectiveness of the proposed method on such datasets with an increasing number of examples.
-Specifically, we consider creating an imbalanced annotated dataset by increasing the number of examples in head classes while keeping the number of examples in tail classes constant.
-For example, in the case of Yahoo (Zhang et al., 2015), when the imbalance ratio $\phi=0$, the number of annotated examples for each class remains 50.
+While our analysis has shown that imbalanced annotations negatively impact ICL by reducing the number of examples in tail classes, one might wonder if this negative effect is due to the reduced size of the dataset rather than a shift in class prior distribution.
+To address this, we verify the negative impact of imbalanced annotations and the effectiveness of the proposed method on such datasets with an increasing number of examples.
+Specifically, we consider creating an imbalanced dataset by increasing the number of examples in the head classes while keeping the number of examples in the tail classes constant.
+For example, in the case of Yahoo [11], when the imbalance ratio $\phi=1$, the number of annotated examples for each class remains 50.
 When the imbalance ratio $\phi=100$, the number of examples in the head classes is 5,000, but the number in the tail classes remains 50.
-We report the test accuracy of Yahoo using OPT-6.7B with six existing selection methods.
+We report the test accuracy of four classification datasets with six existing selection methods.
+Figure 4 (c) shows that imbalanced annotations significantly hurt ICL’s performance when more examples are added to the head classes while keeping the number of examples in the tail classes constant.
+This illustrates that the decreasing trend in ICL’s performance mainly depends on the class prior distribution of the annotated dataset.
+Notably, our method significantly improved the ICL’s performance.
+For instance, for a 100 imbalance ratio, our method outperforms vanilla by a large margin of 5.43, indicating the effectiveness of our method against imbalanced annotation.
+7 Conclusion
+In this paper, we presented a new research direction for understanding and mitigating the effect of imbalanced annotations on in-context learning (ICL).
+Our study reveals that imbalanced annotations significantly degrade the performance of ICL across various tasks, while classical rebalancing methods yield limited improvements.
+We propose RCB, a complementary strategy that estimates conditional bias for reweighting the demonstration score in selection, universally enhancing the performance of ICL with imbalanced annotations.
+To the best of our knowledge, our method is the first to tackle the class-imbalanced issue from the perspective of demonstration selection.
+Extensive experiments demonstrate that RCB can enhance the performance of existing selection methods in ICL under class-imbalanced settings.
+Our approach is easy to use in practice, as it is insensitive to the hyperparameters and does not introduce heavy computational cost.
+Limitations.
+Our methods need to select a balanced subset from imbalanced datasets, which might be impractical if there are few examples in the tail classes.
+It might be an interesting direction to explore how to model the bias in a more data-efficient manner.
+Appendix
 […]
-7 Conclusion
-In this paper, we study the effect of the imbalanced annotated dataset on the performance of in-context learning (ICL).
-Both our theoretical analysis and empirical investigations reveal that imbalanced annotated datasets significantly hurt the performance of ICL.
-Meanwhile classical rebalance methods cannot effectively mitigate the imbalanced datasets.
-To address the issue, we introduce our method, a general strategy that can universally enhance the performance of ICL across various imbalanced ratios.
-Our method is motivated by decomposing the distributional differences between annotated and test datasets into class-wise weights and conditional bias.
-Extensive experiments demonstrate that our method can improve the performance of existing selection methods of ICL across various imbalance ratios.
-Our approach is practical to use, as it is insensitive to the hyperparameters and integrates effortlessly with any LLMs.
-Limitations.
-Our methods need to split a balanced validation dataset from imbalanced annotated datasets, and these methods require that the size of examples in tail classes cannot be too small.
-Furthermore, it might be challenging to estimate multiple parameters for the class-conditional biases using balanced datasets with limited size.
-It might be an interesting direction to explore how to model the bias in a more data-efficient manner.
-Impact Statements
-This paper presents work whose goal is to advance the field of Machine Learning.
-There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
-Appendix A Appendix
-A.1 Related Work
-In-context learning In-context learning (ICL) is a new paradigm for large language (LLMs), which allows LLMs to make predictions only based on a few demonstrations without explicitly updating parameters (Akyurek et al., 2023; Hendel et al., 2023; Agarwal et al., 2024; Dong et al., 2024; Edwards & Camacho-Collados, 2024; Falck et al., 2024).
-Many studies show that ICL can achieve performance similar to fine-tuning but without the high computational cost (Gonen et al., 2023; Mosbach et al., 2023; Müller et al., 2024; Panwar et al., 2024).
-Despite achieving such outstanding performance, ICL has been criticized for being very sensitive to the quality of in-context examples (Fei et al., 2023; Gao et al., 2024a).
-Various approaches have been proposed to improve the robustness of ICL in recent years, including meta-tuning LLMs (Brunet et al., 2023), calibration (Abbas et al., 2024), demonstration selection (Zhang et al., 2022b; Nguyen & Wong, 2023; Qin et al., 2024; Ye et al., 2023b; Gao et al., 2024b; Luo et al., 2024; Mo et al., 2024), ordering (Lu et al., 2022; Liu et al., 2024), number (Zhang et al., 2025) and formation (Voronov et al., 2024; Yao et al., 2024).
-Notably, existing studies often assume that usually assume that annotated dataset and testing data are both i.i.d.
-(independent and identically distributed) sampled from the same data distribution (Luo et al., 2024; Van et al., 2024).
-However, this is not always true for real-world scenarios.
-Specifically, annotated dataset typically exhibit a long-tailed class distribution, where a small portion of classes have a massive number of sample points but the others are associated with only a few examples (Jamal et al., 2020; Schultheis et al., 2024).
-It is still mysterious how demonstrations sampled from long-tailed distribution affect the performance of ICL on both text classification and generation tasks.
-In this paper, we show that annotation quality is crucial for ICL in both text classification and generation, where imbalanced datasetss significantly hurt the performance.
-Different from previous study (Hong et al., 2024), we propose a simple and effective method to enhance the performance of ICL from the perspective of demonstration selection.
-Learning with imbalanced datasets Imbalanced dataset is common in many real-world datasets (Cui et al., 2018, 2019; Jamal et al., 2020; Schultheis et al., 2024).
-The existing approaches to learning with imbalanced datasets can be classified into two types: (1) training imbalance-robust models with imbalance training datasets: designing imbalance-robust loss function (Jamal et al., 2020; Tan et al., 2020; Park et al., 2023; Bhat et al., 2023; Garcin et al., 2022) or designing imbalance-robust model architectures (Long et al., 2022; Pan et al., 2024) to mitigate the issue of imbalanced datasets.
-However, this method is not suitable for ICL, which usually hypothesizes that users are unable to apply fine-tuning techniques.
-(2) handling imbalance examples: handling imbalanced datasets is crucial for ensuring balanced performance across all classes.
-One simple and intuitive approach to deal with the class-imbalanced problem is re-sampling.
-Under-sampling methods remove examples from the majority classes, which is infeasible under data imbalanced settings (Liu et al., 2008; Wei et al., 2022).
-The over-sampling method adds repeated examples for the minority classes, usually causing over-fitting to the minority classes (Chawla et al., 2002; Shi et al., 2023).
-It is noteworthy that these existing methods mainly focus on the fine-tuning setting, and the literature on how to mitigate the effects of imbalanced datasets in in-context learning is limited.
-This motivated us to design a method for addressing the issue of imbalanced datasets under the setting of ICL.
-A.2 Proof
+Appendix A Related Work
+In-context learning In-context learning (ICL) is a new paradigm for large language (LLMs), which allows LLMs to make predictions only based on a few demonstrations without explicitly updating parameters [33, 34, 35, 36, 37, 38].
+Many studies show that ICL can achieve performance similar to fine-tuning but without the high computational cost [39, 2, 40, 3].
+Despite achieving such outstanding performance, ICL has been criticized for being very sensitive to the quality of in-context examples [41, 42].
+Various approaches have been proposed to improve the robustness of ICL in recent years, including meta-tuning LLMs [43], calibration [44], demonstration selection [45, 46, 47, 23, 48, 49, 50], ordering [51, 52], number [53] and formation [54, 55].
+The ICL is closely related to the concept of label bias [1], where language models are biased toward certain answers during few-shot learning.
+In particular, the most relevant part is the majority label bias, a type of context label bias that leads the model to predict answers that appear frequently in the prompt [56, 57].
+An empirical work [6] suggested that label imbalance does not substantially impact ICL performance in binary classification.
+The subsequent works [41, 58, 59, 60] delved deeper into this challenge and proposed to address it by calibrating the model’s output probabilities to compensate for this bias.
+While prior studies have typically focused on imbalanced distributions within prompts, our work shifts the emphasis to class imbalance in the annotation datasets from which demonstrations are drawn.
+Notably, this study represents the first systematic attempt to tackle the class-imbalanced issue from the perspective of demonstration selection.
+Learning with imbalanced datasets Imbalanced datasets are common in many real-world datasets, so the challenge of class imbalance has been widely studied in the literature [61, 9, 28, 62].
+There are two popular directions in learning with imbalanced datasets: (1) Re-sampling: Re-sampling is a widely used strategy in class-imbalanced learning [24], such as Over-sampling [25], which involves repeating data from the minority classes; Under-sampling [26], which involves removing a proportion of data from the majority classes.
+Stratified sampling [27] samples from each class have an identical probability of being sampled.
+In this work, we demonstrate that existing rebalancing methods yield limited improvement in in-context learning (ICL) when dealing with imbalanced annotations.
+(2) Training imbalance-robust models: designing loss function [28, 63, 64, 65, 66] or designing model architectures [67, 68] to mitigate the issue of imbalanced datasets.
+However, this method inevitably incurs high training costs that might be impractical for LLMs.
+In contrast, our approach only estimates the conditional bias for each class, resulting in negligible computational costs compared to traditional training methods.
+Appendix B Proof
 Proof. We apply the importance sampling trick to connect the expected $\operatorname{Error}$ with the imbalanced annotated dataset:
@@
 $=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right]\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{c}(\mathrm{\mathbf{y}})},$
-Here, we decompose the term $\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{c}(\mathrm{\mathbf{y}})}$ into two parts: the class-wise weights $\bm{w}=$, and the conditional weight $\bm{\widetilde{\beta}}=$.
-The class-wise weights measure the difference between class prior distributions $P(Y)$, and the conditional weight measures the difference between conditional distributions $P(X|Y)$.
+Here, we decompose the term $\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{c}(\mathrm{\mathbf{y}})}$ into two parts: the class weights $\bm{w}=P_{t}(\mathrm{\mathbf{y}})/P_{c}(\mathrm{\mathbf{y}})$, and the conditional weight $\bm{\widetilde{\beta}}=\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}$ The class weights measure the difference between class prior distributions $P(Y)$, and the conditional weight measures the difference between conditional distributions $P(X|Y)$.
 Then we have
 $\operatorname{Error}=\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right]\bm{w}(1+\bm{\widetilde{\beta}})$
-where $\bm{w}$ denotes class-wise weights and conditional weight $\bm{\widetilde{\beta}}=\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1$ measures the difference of conditional distribution between the annotated and test datasets.
+where $\bm{w}$ denotes class weights and conditional weight $\bm{\widetilde{\beta}}=\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1$ measures the difference of conditional distribution between the annotated and test datasets.
 For simplicity, we introduce a conditional bias $\bm{\beta}=\bm{w}\times\bm{\widetilde{\beta}}$ and re-write the expected $\operatorname{Error}$ as:
-$\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right](\bm{w}+\bm{\beta}).$ (10)
-where conditional bias $\bm{\beta}=\bm{w}\times\bm{\widetilde{\beta}}=\frac{P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{y}})}\left(\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1\right)$ denotes the deviation of class-wise weights caused by the presence of conditional weight $\bm{\widetilde{\beta}}$.
-A.3 Bayesian Optimizing Conditional Bias
-In this section, we will introduce how to use Bayesian optimization (Gardner et al., 2014; Nogueira, 2014) to find the optimal value of conditional bias $\beta$ by maximizing empirical $\operatorname{Error}$ as follows:
-$\mathcal{F}_{\theta}\left(\beta|\mathcal{D}_{r},\mathcal{D}_{b},\bm{w}\right)=\frac{1}{|\mathcal{D}_{b}|}\sum_{i=1}^{|\mathcal{D}_{b}|}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{(w_{i}+\beta)\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{b})\right\}_{i=1}^{|\mathcal{D}_{r}|}\right),\mathrm{\mathbf{x}}_{b}\right),\mathrm{\mathbf{y}}_{b}\right].$ (11)
+$\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right](\bm{w}+\bm{\beta}).$
+where conditional bias $\bm{\beta}=\bm{w}\times\bm{\widetilde{\beta}}=\frac{P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{y}})}\left(\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1\right)$ denotes the deviation of class weights caused by the presence of conditional weight $\bm{\widetilde{\beta}}$.
+Appendix C Estimating Conditional Bias by Bayesian optimization
+In this section, we will introduce how to use Bayesian optimization [69, 70] to estimate conditional bias $\bm{\beta}$ as follows:
+$\bm{\beta}=\mathop{\arg\min}_{\bm{\beta}}\frac{1}{|\mathcal{D}_{b}|}\sum_{i=1}^{|\mathcal{D}_{b}|}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\{(\bm{w}+\bm{\beta})\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{b})\right\}_{i=1}^{|\mathcal{D}_{r}|}),\mathrm{\mathbf{x}}_{b}\right),\mathrm{\mathbf{y}}_{b}\right],$ (11)
 Surrogate Model.
@@
 Initially, a prior distribution is established in a general form for the estimated model parameters,
-$\mathcal{F}_{\theta}\left(\beta|\mathcal{D}_{r},\mathcal{D}_{b},\bm{w}\right)=\mathcal{N}\left(\mu(\beta),\sigma^{2}(\beta)\right),$
-where $\mathcal{N}$ is the Gaussian distribution with a mean function $\mu(\beta)$ and a covariance function $\sigma^{2}(\beta)$.
-For a data point $\beta_{j}$ sampled from the Gaussian process $\mathcal{N}$, we compute its corresponding function values $\mathcal{F}_{\theta}\left(\beta_{j}|\mathcal{D}_{r},\mathcal{D}_{b},w\right)$.
-Acquisition Function. We use the Expected Improvement (EI) criterion to select the point $\beta_{j+1}$ with the maximum expected improvement as the next query point:
-$\operatorname{EI}(\beta)$ =$=\left\{\begin{array}[]{ll}\left(\mu_{j}(\beta)-\mathcal{F}_{\theta}\left(\beta^{+}|\mathcal{D}_{r},\mathcal{D}_{b},w\right)-\epsilon\right)\Phi(Z)+\sigma_{j}(\beta)\phi(Z)&\text{ if }\sigma_{j}(\beta)>0,\\ 0&\text{ if }\sigma_{j}(\beta)=0.\end{array}\right.$ (13)
-$Z$ $=\frac{\mu_{j}(\beta)-\mathcal{F}_{\theta}\left(\beta^{+}|\mathcal{D}_{r},\mathcal{D}_{b},w\right)-\epsilon}{\sigma_{j}(\beta)}.$
-where $\mu_{j}(\beta)$ and $\sigma_{j}(\beta)$ represent the mean and standard deviation of the surrogate model at step $j$, $\beta^{+}$ denotes the current optimal observed point, $\epsilon$ is a tunable hyper-parameter.
+$\mathcal{F}_{\theta}\left(\bm{\beta}\right)=\mathcal{N}\left(\mu(\bm{\beta}),\sigma^{2}(\bm{\beta})\right),$ (12)
+where $\mathcal{N}$ is the Gaussian distribution with a mean function $\mu(\bm{\beta})$ and a covariance function $\sigma^{2}(\bm{\beta})$.
+For a data point $\bm{\beta}_{j}$ sampled from the Gaussian process $\mathcal{N}$, we compute its corresponding function values $\mathcal{F}_{\theta}\left(\bm{\beta}_{j}\right)$.
+Acquisition Function. We use the Expected Improvement (EI) criterion to select the point $\bm{\beta}_{j+1}$ with the maximum expected improvement as the next query point:
+$\operatorname{EI}(\bm{\beta})$ =$=\left\{\begin{array}[]{ll}\left(\mu_{j}(\bm{\beta})-\mathcal{F}_{\theta}\left(\bm{\beta}^{+}\right)-\epsilon\right)\Phi(Z)+\sigma_{j}(\bm{\beta})\phi(Z)&\text{ if }\sigma_{j}(\bm{\beta})>0,\\ 0&\text{ if }\sigma_{j}(\bm{\beta})=0.\end{array}\right.$ (15)
+$Z$ $=\frac{\mu_{j}(\bm{\beta})-\mathcal{F}_{\theta}\left(\bm{\beta}^{+}\right)-\epsilon}{\sigma_{j}(\bm{\beta})}.$ (16)
+where $\mu_{j}(\bm{\beta})$ and $\sigma_{j}(\bm{\beta})$ represent the mean and standard deviation of the surrogate model at step $j$, $\bm{\beta}^{+}$ denotes the current optimal observed point, $\epsilon$ is a tunable hyper-parameter.
 $\Phi(Z)$ and $\phi(Z)$ are the probability density function and cumulative density function of the Gaussian Process, respectively.
 Iterative Optimization Process.
-We start with an initial set of observations $\beta_{0},...,\beta_{h}$ by evaluating the objective function $\left\{\mathcal{F}\left(\beta_{i}|\mathcal{D}_{r},\mathcal{D}_{b},w,\theta\right)\right\}_{i=0}^{h}$ at a few selected points.
-Second, we use the observed data $\left(\beta_{j},\mathcal{F}\left(\beta_{j}|\mathcal{D}_{r},\mathcal{D}_{b},w,\theta\right)\right)$ to update the Gaussian Process model, refining the estimates of $\mu_{j}(\beta)$ and $\sigma_{j}(\beta)$.
-Third, we employ the Expected Improvement criterion to determine the next point $\beta_{j+1}$ and obtain $\left\{\mathcal{F}\left(\beta_{j+1}|\mathcal{D}_{r},\mathcal{D}_{b},w,\theta\right)\right\}_{i=0}^{h}$.
+We start with an initial set of observations $\bm{\beta}_{0},...,\bm{\beta}_{h}$ by evaluating the objective function $\left\{\mathcal{F}\left(\bm{\beta}_{i}\right)\right\}_{i=0}^{h}$ at a few selected points.
+Second, we use the observed data $\left(\bm{\beta}_{j},\mathcal{F}\left(\bm{\beta}_{j}\right)\right)$ to update the Gaussian Process model, refining the estimates of $\mu_{j}(\bm{\beta})$ and $\sigma_{j}(\bm{\beta})$.
+Third, we employ the Expected Improvement criterion to determine the next point $\bm{\beta}_{j+1}$ and obtain $\left\{\mathcal{F}\left(\bm{\beta}_{j+1}\right)\right\}_{i=0}^{h}$.
 We continue the process of updating the model and selecting new points until a stopping criterion is met.
-Finally, we select the point $\beta^{+}$ with the best observed value of the objective function $\mathcal{F}_{\theta}\left(\beta^{+}|\mathcal{D}_{r},\mathcal{D}_{b},w\right)$ as the optimal solution.
-A.4 Experimental Setting
-Datasets We conduct experiments on various classification and generation tasks and examples of each dataset are shown in Tables 9 and 10.
+Finally, we select the point $\bm{\beta}^{+}$ with the best observed value of the objective function $\mathcal{F}_{\theta}\left(\bm{\beta}^{+}\right)$ as the optimal solution.
+Appendix D Experimental Setting
+D.1 Datasets
+Datasets We conduct experiments on various classification and generation tasks and examples of each dataset are shown in Tables 12 and 13.
 We collect all datasets from Huggingface.
@@
 We randomly subsample examples from the test dataset to verify the performance of the imbalanced annotated dataset.
-We generate categories for Natural question (Kwiatkowski et al., 2019) using ChatGPT-3.5-Turbo.
+We generate categories for Natural question [13] using ChatGPT-3.5-Turbo.
 The categories include people, time, geography, culture, and specialized knowledge.
+D.2 Baselines
 Baselines To verify the effectiveness of our method, we compare our method with previous methods for demonstration retrieval by the downstream ICL’s performance.
@@
   1. 1.
-    Random selects demonstrations randomly from an example set without repetition (Min et al., 2022).
+    Random selects demonstrations randomly from an example set without repetition [30].
   2. 2.
-    TopK retrieves demonstrations that are semantically similar to a test query sample (Liu et al., 2022).
+    TopK retrieves demonstrations that are semantically similar to a test query sample [4].
   3. 3.
-    DPP uses the original BERT embeddings as mentioned above without fine-tuning, and adopts MAP inference for subset retrieval (Ye et al., 2023a).
+    DPP uses the original BERT embeddings as mentioned above without fine-tuning, and adopts MAP inference for subset retrieval [17].
   4. 4.
-    VoteK proposes an unsupervised and graph-based selective annotation method to select diverse and representative demonstrations (Su et al., 2023).
+    VoteK proposes an unsupervised and graph-based selective annotation method to select diverse and representative demonstrations [31].
   5. 5.
-    ConE searches for demonstrations by minimizing the difference in cross-entropy between the test input and the demonstrations (Peng et al., 2024).
+    ConE searches for demonstrations by minimizing the difference in cross-entropy between the test input and the demonstrations [15].
   6. 6.
-    ByCS assumes that an accurate inverse likelihood probability will lead to an accurate posterior probability and selects demonstrations based on their inverse inference results (Wang et al., 2024).
-Experiment details We run our experiments on NVIDIA GeForce RTX 4090 and NVIDIA L40 GPU, and implement all methods by PyTorch and transformers.
-Our code is inspired by OpenICL (Wu et al., 2023).
-We thank the authors for releasing their code.
-A.5 Inference
-Perplexity For classification tasks, we compute the sentence perplexity for each sequence formed by concatenating the input with each candidate answer (Brown et al., 2020; Wu et al., 2023).
+    ByCS assumes that an accurate inverse likelihood probability will lead to an accurate posterior probability and selects demonstrations based on their inverse inference results [6].
+D.3 Inference
+Perplexity For classification tasks, we compute the sentence perplexity for each sequence formed by concatenating the input with each candidate answer [1, 71].
 Specifically, for each input instance $\mathrm{\mathbf{x}}$ is paired potential label set $\mathcal{Y}$, where $\mathcal{Y}$ represents the set of possible classes (e.g., Sports, Business, etc.).
@@
 We select the label corresponding to the input-output pair $\mathrm{\mathbf{c}}$ with the lowest perplexity as the predicted label for the input $\mathrm{\mathbf{x}}$.
-Direct For text generation tasks, we represent candidate answers using tokens from the vocabulary and select the final prediction based on the one with the highest probability (Brown et al., 2020; Wu et al., 2023).
+Direct For text generation tasks, we represent candidate answers using tokens from the vocabulary and select the final prediction based on the one with the highest probability [1, 71].
 To reduce computational cost, we set the maximum new tokens to 50 and limit unnecessary token generation.
-A.6 More empirical results
-A.6.1 The effect of imbalanced dataset
-In this section, we show the effect of imbalanced datasets on the performance of in-context learning with different model architectures and various numbers of subsets.
-Figures LABEL:figure:a1 (b) and (c) report that imbalanced datasets significant degrade ICL’s performance of Llama3-8B (AI@Meta, 2024) and Mistral-7B (Jiang et al., 2023).
-This illustrates that the negative effect of imbalanced datasets can be commonly observed in various large language models.
-Additionally, Figure LABEL:figure:a1 (a) also reports a larger set of demonstrations does not enhance the performance of ICL.
-Specifically, imbalanced datasets also negatively impact the performance of in-context learning.
-A.6.2 The failure of classical rebalance methods
-Figures LABEL:figure:a3 (a) and (b) reports the Kullback-Leibler (KL) divergence between $P_{c}(X|Y)$ and $P_{t}(X|Y)$ for each class in the Yelp and Yahoo datasets, respectively.
-Notably, the KL divergence between $P_{c}(X|Y)$ and $P_{t}(X|Y)$ for each class is unequal to zero and the assumption that $\beta$ = 0 is difficult to be satisfied.
-A.6.3 Main results for each dataset
+D.4 Experiment details
+Experiment details We run our experiments on NVIDIA GeForce RTX 4090 and NVIDIA L40 GPU, and implement all methods by PyTorch and transformers.
+Our code is inspired by OpenICL [71].
+We thank the authors for releasing their code.
+Appendix E More empirical results
+E.1 Performance on different model architectures and sizes
+We conduct experiments on various sized LLMs (including open-weight models and APIs).
+The results (vanilla/ours) in the table 2 below present the same phenomenon as the main experiments in the manuscript: the ICL performance of LLMs get worse at larger imbalanced ratios, and our method can significantly improve the performance, especially at large imbalanced ratios (e.g., 100).
+Table 2: Average accuracy (%) of Vanilla / +Ours methods on the AgNews and Yahoo datasets with different model architectures and sizes. Bold numbers are superior results.
+Imbalance Ratio 1 10 50 100
+Vanilla/+Ours
+OPT-6.7B 68.71/68.63 62.90/67.54 57.35/64.92 54.31/62.07
+OPT-13B 72.22/72.84 68.04/70.12 62.28/68.16 58.12/64.33
+OPT-30B 76.34/76.60 72.32/74.10 67.74/71.90 64.30/70.08
+LLAMA-3-8B 78.90/79.10 75.20/77.16 69.32/75.82 65.72/73.20
+LLAMA-3-70B 86.50/86.43 83.07/84.80 79.86/83.13 77.60/81.22
+ChatGPT-3.5-Turbo 82.45/82.40 79.47/81.08 77.88/80.20 76.32/79.74
+Gemini-2.0-Flash 81.91/82.09 79.95/81.86 77.19/80.31 75.01/79.82
+E.2 Can our method improve real-world imbalanced dataset
+We verify the effectiveness of our method on a real-world imbalanced dataset.
+The Emotion [12] dataset has a long-tailed distribution (refer to Figure 5).
+Table 3 shows that our method consistently improves existing selection methods and outperforms existing rebalancing methods, including over-sampling [25], under-sampling [26], stratified sampling [27], and re-weighting [9].
+For example, with OPT-6.7B [19], using our method boosts the average Macro-F1 of six selection methods from 32.12 to 36.69, a direct improvement of 4.57.
+These results verify that our method is effective in improving ICL’s performance in real-world imbalanced scenarios.
+Figure 5: An example of imbalanced dataset.
+In Emotion [12], a few sentiments make a large contribution and data tend to show a long-tailed distribution.
+For example, “joy“ and “sadness“ are head classes, while most other classes, such as “love“ and “surprise“, are tail classes.
+Table 3: Average Macro-F1 metric on Emotion dataset. Bold numbers are superior results.
+Methods Vanilla +Over-sampling +Under-sampling +Re-Weighting +Stratified-Sampling +Ours
+Accuracy 32.22 33.74 34.22 30.27 31.60 36.69
+E.3 Reweighting with conditional bias vs. output calibration?
+The ICL is closely related to the concept of label bias [1], where language models are biased toward certain answers during few-shot learning.
+The subsequent works [41, 58, 59, 60] delved deeper into this challenge and proposed to address it by calibrating the model’s output probabilities to compensate for this bias.
+One may also ask: can a similar effect be achieved by calibrating the output of ICL with imbalanced demonstrations?
+In this ablation, we compare our method with some calibration methods, including CC [56], DC [41] and LOOC [60].
+Table 4 presents that although both the demonstration selection and calibration methods improve the performance of ICL against imbalanced annotations, the calibration methods obtain inferior performance compared to our proposed method.
+This is because the calibration methods [56, 41, 60] only ensure that ICL with imbalanced demonstrations generates consistent output with that of balanced demonstrations.
+However, some effective selection methods, such as TopK [4] and DPP [23], are more likely to select examples belonging to the same class as the test sample for demonstrations and conduct imbalanced demonstrations.
+For these selection methods, the calibration methods inevitably degrade the performance of ICL, whereas our method can preserve their effectiveness.
 […]
-A.6.4 Is our method effective with different model architectures?
-In this section, we provide the detailed results of Llama3-8B (AI@Meta, 2024) and Mistral-7B (Jiang et al., 2023).
-Figures LABEL:figure:a4 (a) and (b) report the average accuracy of four classification tasks across six selection methods.
-We can clearly observe that imbalanced annotated datasets significantly hurt the performance of in-context learning, which is consistent with our previous findings.
-Meanwhile, Figures LABEL:figure:a4 (a) and (b) also demonstrate that integrating our method can improve the performance of ICL against imbalanced annotated datasets.
-A.6.5 Can our method improve real-world imbalanced dataset?
-We further verify the effectiveness of our method on a real-world class-imbalanced dataset (Emotion (Saravia et al., 2018)).
-.We select demonstrations from the training set and construct a balanced test set from the original test set.
-Figure 18 summarizes the test accuracy across six existing selection methods using baselines, classical rebalance methods, and our method on the Emotion dataset.
-In particular, the proposed method is able to consistently improve the existing selection methods and outperforms existing rebalance methods in test accuracy.
+Imbalance Ratios 1 10 50 100
+Vanilla 84.45 75.19 67.78 64.02
++CC 78.38 77.29 75.75 74.42
++DD 79.63 78.67 76.71 75.58
++LOOC 75.71 73.00 74.25 73.88
++Ours 84.26 83.35 80.28 75.28
+E.4 Discussion of imbalanced test set
+We also conducted experiments on an imbalanced test set.
+The table below presents the average Macro-F1 metric on four datasets: AgNews, Amazon, Yelp, and Yahoo.
+The results in Table 5 demonstrate that our method can improve ICL performance on an imbalanced test set.
+For example, on the AgNews and Yahoo datasets with 100 imbalance ratio, our approach improves the test Macro-F1 of vanilla ICL from 45.42 to 53.85 – a 8.43 of direct improvement.
+Table 5: Average Macro-F1 metric of Vanilla / +Ours methods on the AgNews and Yahoo datasets with imbalanced test set. Bold numbers are superior results.
+Imbalance Ratios 1 10 50 100
+Vanilla 64.32 58.11 54.09 45.42
++Ours 64.52 61.67 57.89 53.85
+E.5 Results of extreme imbalanced ratios
+[…]
+Table 6: Average accuracy (%) of Vanilla / +Ours methods on the AgNews and Yahoo datasets with extreme imbalance ratios. Bold numbers are superior results.
+Imbalance Ratios 10 100 500 1000
+Vanilla 62.90 54.31 46.80 44.52
++Ours 67.54 62.07 58.34 56.33
+Improvement 4.64 7.76 11.54 11.81
+E.6 The computational cost of our method
+In this section, we illustrate the high efficiency of our method from the following aspects.
+Firstly, our method estimates the conditional bias by accessing low-cost model outputs and does not require extensive hyperparameter tuning.
+The table 7 reports the average computational cost (in minutes) between our method and parameter-efficient fine-tuning (PEFT) [72] on the AgNews dataset across various imbalanced ratios.
+For example, with an imbalanced ratio of 100, one iteration of Bayesian optimization can be completed in 2.51 minutes using OPT-6.7B on a single NVIDIA L40 GPU.
+Besides, the bayesian optimization employed in our method is well-suited for efficiently optimizing non-differentiable and black-box functions with a few iterations [69, 73, 74].
+In fact, our method can achieve significant improvement with minimal computational cost using 30 iterations and a balanced subset of size 100.
+Our method is acceptable in practice and only model weights $w+\beta$ one time for the whole test dataset.
+The detailed settings of PEFT are as follows: Learning Rate (1e-5); LR Scheduler Type (cosine); Batch Size (8); Number of Training Epochs (5); Number of Warmup Steps (100); Weight Decay (0.01); Lora Alpha (32); Lora Dropout (0.05).
+Table 7: Comparison of average computational cost (minute) per iteration using RCB and PEFT on AgNews. Bold numbers are superior results.
+Imbalance Ratios 1 10 50 100
+PEFT 72.35 35.25 27.52 24.38
+RCB 3.04 2.58 2.53 2.51
+E.7 Main results for each dataset
+[…]
 […]
 $K$ 8 16
-Imbalance Ratio 0 10 50 100 0 10 50 100
+Imbalance Ratio 1 10 50 100 1 10 50 100
 Random 68.17$\pm$0.12 58.58$\pm$1.25 48.75$\pm$1.81 46.25$\pm$0.35 70.58$\pm$0.59 61.33$\pm$0.82 50.75$\pm$1.62 48.25$\pm$1.24
@@
 +Ours 84.73$\pm$0.38 82.75$\pm$1.24 80.92$\pm$1.36 77.17$\pm$1.85 88.33$\pm$0.12 86.25$\pm$1.41 81.50$\pm$1.41 72.75$\pm$1.41
-Baselines 82.92$\pm$0.32 73.03$\pm$1.44 65.81$\pm$1.31 62.45$\pm$1.31 84.45$\pm$0.39 75.19$\pm$0.81 67.78$\pm$1.21 64.02$\pm$0.70
+Vanilla 82.92$\pm$0.32 73.03$\pm$1.44 65.81$\pm$1.31 62.45$\pm$1.31 84.45$\pm$0.39 75.19$\pm$0.81 67.78$\pm$1.21 64.02$\pm$0.70
 +Ours 82.75$\pm$0.59 82.20$\pm$1.05 79.24$\pm$1.27 75.67$\pm$1.69 84.26$\pm$0.32 83.35$\pm$1.21 80.28$\pm$1.20 75.28$\pm$1.95
@@
 $K$ 8 16
-Imbalance Ratio 0 10 50 100 0 10 50 100
+Imbalance Ratio 1 10 50 100 1 10 50 100
 Random 37.93$\pm$1.91 39.53$\pm$2.22 36.07$\pm$0.41 33.13$\pm$1.47 44.20$\pm$0.59 44.47$\pm$0.81 39.27$\pm$1.18 37.60$\pm$0.59
@@
 +Ours 53.40$\pm$0.43 52.33$\pm$0.90 48.47$\pm$1.65 46.50$\pm$0.57 55.37$\pm$0.68 52.07$\pm$0.94 50.40$\pm$0.86 48.87$\pm$0.94
-Baselines 50.34$\pm$0.56 48.90$\pm$0.84 44.20$\pm$1.07 41.50$\pm$1.64 52.96$\pm$0.48 50.61$\pm$0.93 46.91$\pm$0.88 44.60$\pm$1.12
+Vanilla 50.34$\pm$0.56 48.90$\pm$0.84 44.20$\pm$1.07 41.50$\pm$1.64 52.96$\pm$0.48 50.61$\pm$0.93 46.91$\pm$0.88 44.60$\pm$1.12
 +Ours 50.32$\pm$0.70 49.26$\pm$0.73 45.96$\pm$1.36 45.53$\pm$1.64 53.00$\pm$0.64 51.72$\pm$0.90 49.56$\pm$1.20 48.86$\pm$1.53
@@
 $K$ 8 16
-Imbalance Ratio 0 10 50 100 0 10 50 100
-Random 12.33$\pm$0.62 15.2$\pm$1.45 18.53$\pm$1.32 20.13$\pm$0.98 13.33$\pm$0.19 18.53$\pm$1.05 20.07$\pm$0.09 20.33$\pm$0.34
-+Ours 12.53$\pm$0.47 19.87$\pm$0.96 20.93$\pm$2.38 20.47$\pm$1.33 13.47$\pm$0.25 16.13$\pm$2.54 20.20$\pm$1.18 21.07$\pm$0.62
-TopK 28.73$\pm$0.09 28.47$\pm$0.74 24.40$\pm$1.80 23.40$\pm$0.86 30.60$\pm$0.00 28.13$\pm$1.15 24.20$\pm$0.65 23.07$\pm$0.66
-+Ours 28.80$\pm$0.00 29.47$\pm$0.50 27.87$\pm$0.96 27.20$\pm$0.49 30.40$\pm$0.00 31.00$\pm$1.30 29.13$\pm$0.93 28.33$\pm$0.90
-DPP 29.80$\pm$0.16 27.13$\pm$0.68 25.47$\pm$1.27 24.13$\pm$0.5 32.47$\pm$0.09 28.13$\pm$0.47 24.00$\pm$0.43 23.00$\pm$0.75
-+Ours 29.60$\pm$0.00 28.67$\pm$1.79 28.67$\pm$0.81 27.60$\pm$0.86 31.87$\pm$0.75 30.87$\pm$1.39 28.93$\pm$1.00 27.93$\pm$1.06
-VoteK 28.67$\pm$0.09 25.67$\pm$0.81 22.73$\pm$1.23 22.93$\pm$0.41 31.57$\pm$0.29 27.67$\pm$0.82 22.60$\pm$0.65 21.93$\pm$0.25
-+Ours 28.13$\pm$0.09 27.53$\pm$1.72 27.00$\pm$0.75 27.47$\pm$1.32 31.80$\pm$0.85 30.07$\pm$0.34 28.40$\pm$1.40 27.20$\pm$1.02
-ConE 26.73$\pm$0.47 25.40$\pm$1.34 22.87$\pm$0.82 21.27$\pm$1.16 30.20$\pm$0.00 27.73$\pm$0.47 22.87$\pm$0.94 22.27$\pm$0.38
-+Ours 27.40$\pm$0.82 26.73$\pm$1.27 27.60$\pm$1.30 27.27$\pm$1.16 30.27$\pm$0.34 29.07$\pm$1.20 29.13$\pm$0.93 26.80$\pm$0.49
-ByDC 29.47$\pm$0.52 27.13$\pm$0.34 25.20$\pm$1.02 23.53$\pm$1.57 32.13$\pm$0.38 27.93$\pm$0.66 24.13$\pm$0.50 22.27$\pm$0.25
-+Ours 29.67$\pm$0.50 28.80$\pm$1.42 29.20$\pm$0.71 27.13$\pm$1.05 31.53$\pm$1.23 30.07$\pm$1.31 28.60$\pm$1.14 27.8$\pm$0.75
-Baselines 25.96$\pm$0.33 24.83$\pm$0.89 23.20$\pm$1.24 22.57$\pm$0.91 28.38$\pm$0.16 26.35$\pm$0.77 22.98$\pm$0.54 22.15$\pm$0.44
-+Ours 26.02$\pm$0.31 26.85$\pm$1.28 26.88$\pm$1.15 26.19$\pm$1.04 28.26$\pm$0.57 27.87$\pm$1.35 27.40$\pm$1.10 26.52$\pm$0.81
+Imbalance Ratio 1 10 50 100 1 10 50 100
+Random 42.07$\pm$0.69 36.97$\pm$0.64 35.93$\pm$0.71 35.47$\pm$0.38 43.67$\pm$0.44 38.70$\pm$0.60 36.27$\pm$0.91 35.57$\pm$0.09
++Ours 41.63$\pm$0.89 41.67$\pm$1.96 38.90$\pm$1.00 40.17$\pm$0.98 43.00$\pm$0.60 43.60$\pm$1.47 42.93$\pm$0.56 43.33$\pm$1.24
+TopK 46.50$\pm$0.00 42.43$\pm$1.02 37.17$\pm$1.36 37.13$\pm$1.31 47.60$\pm$0.00 43.20$\pm$0.27 38.07$\pm$1.29 37.17$\pm$1.36
++Ours 46.13$\pm$0.49 45.27$\pm$0.38 42.47$\pm$1.29 41.23$\pm$0.36 47.30$\pm$0.40 46.57$\pm$0.58 43.60$\pm$0.27 40.87$\pm$1.22
+DPP 46.70$\pm$0.00 41.53$\pm$0.36 38.30$\pm$0.20 37.73$\pm$0.64 48.10$\pm$0.00 43.37$\pm$0.18 39.27$\pm$0.56 37.77$\pm$0.58
++Ours 46.30$\pm$0.27 43.63$\pm$1.11 41.23$\pm$1.96 39.77$\pm$0.56 47.97$\pm$0.09 45.47$\pm$0.44 42.13$\pm$0.96 40.23$\pm$0.44
+VoteK 46.87$\pm$0.05 40.90$\pm$0.64 37.00$\pm$1.39 37.03$\pm$1.56 48.03$\pm$0.04 43.50$\pm$0.53 37.57$\pm$0.62 36.73$\pm$0.98
++Ours 46.67$\pm$0.16 43.23$\pm$1.49 42.50$\pm$1.00 40.57$\pm$0.64 47.93$\pm$0.16 44.83$\pm$0.58 43.17$\pm$0.69 40.73$\pm$0.64
+ConE 46.13$\pm$0.38 40.27$\pm$0.66 36.60$\pm$0.86 35.67$\pm$1.67 48.10$\pm$0.07 42.77$\pm$0.16 37.87$\pm$0.76 36.33$\pm$0.91
++Ours 46.33$\pm$0.44 41.60$\pm$1.20 40.80$\pm$0.93 39.23$\pm$0.91 48.00$\pm$0.13 44.57$\pm$0.84 42.93$\pm$0.62 39.87$\pm$0.42
+ByDC 46.73$\pm$0.09 41.60$\pm$1.40 37.47$\pm$0.93 36.57$\pm$1.10 48.23$\pm$0.09 43.43$\pm$0.11 38.60$\pm$1.00 37.43$\pm$0.71
++Ours 47.07$\pm$0.49 44.27$\pm$0.58 41.37$\pm$1.58 40.03$\pm$1.09 48.03$\pm$0.24 45.47$\pm$0.96 42.67$\pm$0.24 40.57$\pm$0.51
+Vanilla 45.83$\pm$0.20 40.62$\pm$0.79 37.08$\pm$0.91 36.60$\pm$1.11 47.29$\pm$0.11 42.49$\pm$0.31 37.94$\pm$0.86 36.83$\pm$0.77
++Ours 45.69$\pm$0.46 43.28$\pm$1.12 41.21$\pm$1.29 40.17$\pm$0.76 47.04$\pm$0.27 45.08$\pm$0.81 42.91$\pm$0.56 40.93$\pm$0.75
 […]
 $K$ 8 16
-Imbalance Ratio 0 10 50 100 0 10 50 100
-Random 12.07$\pm$0.25 15.73$\pm$0.34 20.27$\pm$1.25 21.33$\pm$0.52 12.87$\pm$0.47 20.33$\pm$0.96 22.87$\pm$2.07 23.87$\pm$0.41
-+Ours 11.27$\pm$0.66 20.13$\pm$0.98 19.13$\pm$1.11 22.40$\pm$2.41 13.20$\pm$0.28 15.60$\pm$0.86 20.40$\pm$1.14 22.80$\pm$0.43
-TopK 31.40$\pm$0.00 29.67$\pm$1.20 25.87$\pm$0.68 25.33$\pm$0.66 36.20$\pm$0.16 32.47$\pm$1.06 27.40$\pm$0.99 25.47$\pm$0.09
-+Ours 30.80$\pm$0.00 29.83$\pm$1.01 27.87$\pm$1.05 26.20$\pm$1.23 33.60$\pm$0.85 31.17$\pm$0.29 29.93$\pm$0.52 27.13$\pm$0.98
-DPP 31.93$\pm$0.09 30.00$\pm$0.49 27.47$\pm$1.04 26.20$\pm$1.02 38.33$\pm$0.25 32.93$\pm$1.11 28.07$\pm$0.93 26.53$\pm$0.25
-+Ours 31.60$\pm$0.85 30.40$\pm$1.57 27.93$\pm$0.93 27.07$\pm$2.10 36.00$\pm$0.00 30.73$\pm$0.66 29.80$\pm$0.71 30.03$\pm$0.73
-VoteK 28.80$\pm$0.16 29.60$\pm$0.59 27.07$\pm$1.39 25.73$\pm$0.41 36.13$\pm$0.09 32.47$\pm$1.06 27.40$\pm$0.99 25.47$\pm$0.09
-+Ours 28.93$\pm$0.74 29.13$\pm$1.48 27.87$\pm$0.62 26.47$\pm$0.66 36.07$\pm$0.34 32.47$\pm$1.06 29.27$\pm$0.57 28.33$\pm$0.82
-ConE 28.47$\pm$1.32 27.60$\pm$1.66 24.93$\pm$1.05 23.87$\pm$0.93 36.87$\pm$0.38 29.93$\pm$0.52 25.93$\pm$0.09 24.13$\pm$0.47
-+Ours 29.13$\pm$0.38 27.93$\pm$0.41 26.33$\pm$1.72 25.20$\pm$0.49 34.53$\pm$0.09 31.07$\pm$0.19 27.60$\pm$1.23 27.47$\pm$0.47
-ByDC 31.73$\pm$0.09 29.27$\pm$0.84 26.80$\pm$1.02 26.33$\pm$0.84 37.53$\pm$0.34 30.87$\pm$0.62 26.47$\pm$0.57 24.80$\pm$0.43
-+Ours 31.47$\pm$0.25 30.27$\pm$0.19 27.80$\pm$0.72 27.27$\pm$0.74 35.20$\pm$0.16 31.53$\pm$0.34 28.80$\pm$1.50 28.13$\pm$1.46
-Baselines 27.40$\pm$0.32 26.98$\pm$0.85 25.40$\pm$1.07 24.80$\pm$0.73 32.99$\pm$0.28 29.83$\pm$0.89 26.36$\pm$0.94 25.05$\pm$0.29
-+Ours 27.20$\pm$0.48 27.95$\pm$0.94 26.16$\pm$1.02 25.77$\pm$1.27 31.43$\pm$0.29 28.76$\pm$0.57 27.63$\pm$0.95 27.32$\pm$0.82
-Table 8: The average Rouge score of CodeSearchNet dataset across six selection methods with various imbalance ratios. Bold numbers are the best results.
-Imbalance Ratio 0 10 50 100
-Baselines 15.99 13.26 11.41 9.84
-+Over-Sampling 16.12 13.82 10.07 8.56
-+Under-Sampling 16.08 10.17 10.06 9.20
-+Stratified-Sampling 10.59 9.95 9.54 9.24
-+Re-Weighting 15.86 13.33 8.59 7.96
-+Ours 15.71 13.64 11.86 10.90
-Figure 18: Average accuracy of Emotion tasks across six existing selection methods using OPT-6.7B (Zhang et al., 2022a).
+Imbalance Ratio 1 10 50 100 1 10 50 100
+Random 42.63$\pm$0.38 39.80$\pm$0.93 37.60$\pm$0.73 37.20$\pm$0.13 44.27$\pm$0.36 42.10$\pm$0.67 38.90$\pm$0.27 38.17$\pm$0.22
++Ours 42.90$\pm$0.80 42.97$\pm$1.42 41.40$\pm$1.07 42.20$\pm$0.13 45.40$\pm$0.60 45.63$\pm$0.36 44.40$\pm$1.73 43.30$\pm$1.40
+TopK 47.30$\pm$0.00 45.73$\pm$0.58 41.13$\pm$0.64 40.27$\pm$1.36 49.70$\pm$0.00 46.97$\pm$0.38 43.30$\pm$0.27 42.17$\pm$0.49
++Ours 46.97$\pm$0.44 45.40$\pm$0.67 43.93$\pm$0.89 42.07$\pm$0.64 48.10$\pm$0.00 46.60$\pm$0.40 44.93$\pm$0.11 43.83$\pm$0.36
+DPP 46.80$\pm$0.00 44.83$\pm$0.82 41.97$\pm$0.84 40.57$\pm$0.44 48.10$\pm$0.00 46.33$\pm$0.89 43.90$\pm$0.13 42.77$\pm$0.18
++Ours 46.27$\pm$0.71 44.60$\pm$0.33 42.37$\pm$0.91 41.97$\pm$0.36 48.13$\pm$0.04 45.83$\pm$0.78 44.20$\pm$0.47 44.03$\pm$0.49
+VoteK 48.10$\pm$0.20 45.60$\pm$0.33 41.77$\pm$0.36 40.73$\pm$0.49 49.20$\pm$0.00 46.70$\pm$0.47 43.40$\pm$0.53 42.07$\pm$0.56
++Ours 47.77$\pm$0.31 45.40$\pm$0.53 42.60$\pm$0.80 41.63$\pm$0.31 49.00$\pm$0.13 46.73$\pm$0.16 44.10$\pm$0.53 43.17$\pm$0.58
+ConE 46.20$\pm$0.00 44.17$\pm$0.69 41.23$\pm$1.18 39.70$\pm$0.47 48.63$\pm$0.04 46.40$\pm$0.53 42.73$\pm$0.56 41.87$\pm$0.29
++Ours 46.53$\pm$0.44 44.50$\pm$0.73 42.57$\pm$0.58 40.87$\pm$0.64 48.43$\pm$0.11 46.57$\pm$0.42 43.23$\pm$0.22 42.53$\pm$0.38
+ByDC 47.70$\pm$0.33 45.27$\pm$0.16 41.93$\pm$0.58 40.80$\pm$0.53 48.97$\pm$0.09 46.37$\pm$0.36 43.07$\pm$0.11 42.00$\pm$0.33
++Ours 47.43$\pm$0.18 45.60$\pm$0.40 42.27$\pm$0.51 42.10$\pm$0.20 48.37$\pm$0.09 46.33$\pm$0.24 43.57$\pm$0.29 43.57$\pm$0.29
+Vanilla 46.46$\pm$0.15 44.23$\pm$0.59 40.94$\pm$0.72 39.88$\pm$0.57 48.14$\pm$0.08 45.81$\pm$0.55 42.55$\pm$0.31 41.51$\pm$0.34
++Ours 46.31$\pm$0.48 44.74$\pm$0.68 42.52$\pm$0.79 41.81$\pm$0.38 47.91$\pm$0.16 46.28$\pm$0.39 44.07$\pm$0.56 43.46$\pm$0.58
 Table 12: The statistics, split and evaluation metrics of each dataset.
@@
 Yahoo 50000 500 10 Accuracy
-Emotion 20000 300 6 Accuracy
+Emotion 15758 1974 6 Macro-F1
 Natural Question 25000 500 5 Exact Match
@@
 […]
+Figure 6: The Kullback-Leibler (KL) divergence between $P_{c}(X|Y)$ and $P_{t}(X|Y)$ for each class in the Yahoo datasets.
+The KL divergence validates the existence of conditional bias in imbalanced datasets.
+(a) Amazon
+(b) Yelp
+(c) Yahoo
+Figure 7: The comparison of average accuracy for each class in the Amazon (a), Yelp (b) and Yahoo (c) datasets across various imbalance ratios.